### v1

# Estimates for Dirichlet Eigenvalues of the Schrödinger operator with the Kronig-Penney Model

Abstract
In this paper, first, we improve the asymptotic formulas obtained in [13] and obtain sharp asymptotic formulas explicitly expressed by the potential. For the potentials of bounded variation, we obtain asymptotic formulas in which the first and second terms are explicitly determined and separated from the error terms. In addition, we illustrate these formulas for the Kronig-Penney potential. We then provide estimates for the small Dirichlet eigenvalues of the one-dimensional Schrödinger operator in the Kronig-Penney model. Using Rouché’s theorem, we derive several useful equations from certain iteration formulas for computing these Dirichlet eigenvalues, and we estimate the eigenvalues numerically. Moreover, we present error estimates and include a numerical example.
Key Words: Eigenvalue estimations, Dirichlet boundary conditions, Kronig-Penney model.
AMS Mathematics Subject Classification: 34L05, 34L15, 65L15.
1 Introduction and Preliminary Facts
This paper is a continuation and contains some applications of the paper [13]. In [13], we considered the operator generated in $L_{2}[0,1]$ by the expression
$y^{\prime\prime}+q(x)y,$ (1)
and the boundary conditions
$y(1)=y(0)=0,$
where $q(x)$ is a complex-valued summable function. In the present paper, for brevity of some calculation, $L_{2}[0,1]$ is replaced by $L_{2}[0,\pi]$ and it is assumed that ${\displaystyle\int_{0}^{\pi}}q(x)dx=0$. Therefore, we consider the operator $L(q)$ generated in $L_{2}[0,\pi]$ by the expression (1) and the boundary conditions
$y(\pi)=y(0)=0,$
where $q(x)$ is a complex-valued summable function.
In [13], we proved that the eigenvalue $\lambda_{n}$ of the operator $L(q)$ lying in $O(1)$ neighborhood of $n^{2}$ satisfies the equality
$(\lambda_{n}-n^{2}-A_{m}(\lambda_{n}))(\Psi_{n}(x),\sin nx)=R_{m+1},$ (2)
where $n$ is a large number, $\Psi_{n}(x)$ is an eigenfunction corresponding to the eigenvalue $\lambda_{n}$ and satisfying the inequality $|(\Psi_{n},\sin nx)|>1/2$,
$A_{m}(\lambda_{n})=\sum_{k=0}^{m}a_{k}(\lambda_{n}),$
$a_{0}(\lambda_{n})=-C_{2n},\qquad a_{1}(\lambda_{n})=\sum_{\begin{subarray}{c}n_{1}=-\infty\\ n_{1}\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{n_{1}}(C_{n_{1}}-C_{n_{1}+2n})}{\lambda_{n}-(n+n_{1})^{2}},$ (3)
$a_{2}(\lambda_{n})=\sum_{\begin{subarray}{c}n_{1},n_{2}=-\infty\\ n_{1},n_{1}+n_{2}\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{n_{1}}C_{n_{2}}(C_{n_{1}+n_{2}}-C_{n_{1}+n_{2}+2n})}{[\lambda_{n}-(n+n_{1})^{2}][\lambda_{n}-(n+n_{1}+n_{2})^{2}]},$ (4)
$a_{k}(\lambda_{n})=\sum_{n_{1},n_{2},...,n_{k}=-\infty}^{\infty}\frac{C_{n_{1}}C_{n_{2}}...C_{n_{k}}(C_{n_{1}+n_{2}+..+n_{k}}-C_{n_{1}+n_{2}+...+n_{k}+2n})}{\prod_{j=1}^{k}[\lambda_{n}-(n+\sum_{s=1}^{j}n_{s})^{2}]},$ (5)
$R_{m+1}(\lambda_{n})=\sum_{n_{1},n_{2},...,n_{m+1}=-\infty}^{\infty}\frac{C_{n_{1}}C_{n_{2}}...C_{n_{m+1}}\bigl(q(x)\Psi_{n}(x),\sin\bigl(n+\sum_{k=1}^{m+1}n_{k}\bigr)\bigr)}{\prod_{j=1}^{m+1}[\lambda_{n}-(n+\sum_{k=1}^{j}n_{k})^{2}]},$ (6)
and $C_{n}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos nxdx$. Here the sums for $a_{k}(\lambda_{n})$ and $R_{m+1}(\lambda_{n})$ are taken under the conditions
 $\text{ }\sum\limits_{j=1}^{s}n_{j}\neq 0,-2n,$
for s=1,2,,ks=1,2,...,k and s=1,2,,m+1s=1,2,...,m+1, respectively. Moreover, we have
$a_{k}=O\biggl(\bigl(\frac{ln\left|n\right|}{n}\bigr)^{k}\biggr),\qquad R_{m+1}=O\biggl(\bigl(\frac{ln\left|n\right|}{n}\bigr)^{m+1}\biggr).$ (7)
Using (2)-(7), we obtained the following asymptotic formulas:
$\lambda_{n}=n^{2}-C_{2n}+O\bigl(\frac{ln\left|n\right|}{n}\bigr),$
$\lambda_{n}=n^{2}-C_{2n}+\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}(C_{k}-C_{k+2n})}{[n^{2}-(n+k)^{2}]}+O\biggl(\bigl(\frac{ln\left|n\right|}{n}\bigr)^{2}\biggr),$ (8)
and other formulas of order $O((\ln n)^{l}n^{-l})$ ( for all $l>0$ ) for $\lambda_{n}$ and corresponding eigenfunctions of the operator $L(q)$.
In this paper, in Section 2, we improve the asymptotic formulas, instead of the series in (8), we write some Fourier coefficients of some functions explicitly expressed by the potential $q$. For the potential of bounded variation, we explicitly determine the first and second terms of the asymptotic formulas, which have the order $O(n^{-1})$ and $O(n^{-2})$, respectively (see (33) and Theorem 1). Note that, generally speaking, if the potential $q$ is a piecewise continuous function with some jump, then the first term $C_{2n}$ is of order $n^{-1}$ and the next term in the asymptotic formula (33) is of order $n^{-2}$. Consequently, asymptotic formula (33) with error $O(n^{-3})$ explicitly separates the first and second terms from errors. Then, we demonstrate this formula for the Kronig-Penney potential
$q(x)=\begin{cases}a&\text{if }x\in[0,c]\\[5.69054pt] b&\text{if }x\in(c,\pi],\end{cases}$ (9)
where $q(x+\pi)=q(x)$, and $c\in(0,\pi)$. Without loss of generality, we assume that $a<b$, and
$q_{0}=\frac{1}{\pi}\int_{0}^{\pi}q(x)dx=0.$
Therefore, we have
$ac+b(\pi-c)=0,\qquad a<0<b,$ (10)
and $(b-a)c=b\pi$.
Note that, the Kronig-Penney model is a simplified model for an electron in a one-dimensional periodic potential and has been studied in many works (see, for example, [1, 3, 8, 10, 11], and references therein). Veliev [10, 11] studied the bands and gaps in the spectrum of the Schrödinger operator with Kronig-Penney potential and obtained asymptotic formulas for the length of the gaps in the spectrum.
In Section 3, we provide estimates for small Dirichlet eigenvalues of the Schrödinger operator $L(q)$ with the Kronig-Penney potential. We obtain some useful equations for calculating the Dirichlet eigenvalues using Rouche’s theorem. These equations are derived from some iterations formulas by the methods used in [7, 9, 12, 13]. It is important to note that in these papers, the authors used asymptotic formulas for large eigenvalues, which cannot be applied to small eigenvalues. Therefore, in Section 3, we examine the small eigenvalues using numerical methods. Moreover, as a result of meticulous and detailed investigations, we establish conditions on the potential under which all small eigenvalues satisfy the equations obtained from the iterative formulas.
2 On the Sharp Asymptotic Formulas
In this section, first, to obtain explicit formulas for the large eigenvalues from (8), we consider the series in (8) as the difference of the series
$A_{1,n}(n^{2}):=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}^{2}}{n^{2}-(n+k)^{2}}$ (11)
and
$B_{1,n}(n^{2})=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}C_{k+2n}}{n^{2}-(n+k)^{2}}.$ (12)
Then, we consider these series in the following lemmas. First, let us consider $A_{1,n}(n^{2})$. For this introduce the notations:
$Q(x,n)=\int_{0}^{x}q(t)e^{i2nt}\,dt-q_{-2n}x,\qquad Q_{n,0}=\frac{1}{\pi}\int_{0}^{\pi}Q(x,n)dx$ (13)
and
$G(x,n)=\int_{0}^{x}\bigl(q(t)e^{i2nt}\,+\frac{i\pi}{2}e^{it}\,q_{-2n}\bigr)dt,$ (14)
where
$q_{k}=\frac{1}{\pi}\int_{0}^{\pi}q(x)e^{-ikx}dx\,,\qquad q_{0}=0.$
Lemma 1
If $q\in L_{1}[0,\pi]$, then the following formula holds:
$A_{1,n}(n^{2})=\frac{C_{2n}^{2}}{4n^{2}}+D(n,q)+2\operatorname{Re}D(n,p),$ (15)
where $D(n,f)=D_{1}(n,f)+D_{2}(n,f)$,
$D_{1}(n,f)=:\frac{i}{4n\pi}\int_{0}^{\pi}f(x)(Q(x,n)-Q_{n,0})e^{-i2nx}\,dx,\text{ }$
$D_{2}(n,f)=:\frac{i}{4n\pi}\int_{0}^{\pi}f(x)G(x,n)e^{-i2nx}\,dx,$
and p(x)=q(x)p(x)=q(-x).
Proof. By the definition of $A_{1,n}(n^{2})$, we have
$A_{1,n}(n^{2})=-\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}^{2}}{k(2n+k)}.$
Since
$\frac{1}{k(2n+k)}=\frac{1}{2n}\biggl(\frac{1}{k}-\frac{1}{2n+k}\biggr),$
we see that
$A_{1,n}(n^{2})=\frac{1}{2n}\sum_{k\neq 0,-2n}\frac{C_{k}^{2}}{2n+k}-\frac{1}{2n}\sum_{k\neq 0,-2n}\frac{C_{k}^{2}}{k}.$ (16)
Using the obvious equality $C_{-k}=C_{k}$ and $C_{0}=0,$ we obtain the following equality for the second term on the right side of (16):
$\frac{1}{2n}\sum_{k\neq 0,-2n}\frac{C_{k}^{2}}{k}=\frac{C_{2n}^{2}}{4n^{2}}.$ (17)
Now let us estimate the first term on the right side of (16). Since $q_{-k}=\overline{q_{k}}$ and $\cos kx=\frac{1}{2}(e^{ikx}+e^{-ikx})$ we have
$C_{k}^{2}=\frac{1}{4}(2\left|q_{k}\right|^{2}+q_{k}^{2}+q_{-k}^{2}).$
Therefore, the first term on the right side of (16) can be written in the form
$\frac{1}{2n}\sum_{k\neq 0,-2n}\frac{C_{k}^{2}}{2n+k}=I_{1}+I_{2}+I_{3},$ (18)
where
$I_{1}=\frac{1}{4n}\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{\left|q_{k}\right|^{2}}{(2n+k)}=I_{1,1}+I_{1,2},\qquad I_{2}=\frac{1}{8n}\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{q_{-k}^{2}}{(2n+k)},\qquad I_{3}=\overline{I_{2}},$ (19)
$I_{1,1}=\frac{1}{4n}\sum_{\begin{subarray}{c}m=-\infty\\ m\neq 0,-n\end{subarray}}^{\infty}\frac{\left|q_{2m}\right|^{2}}{(2n+2m)},\qquad I_{1,2}=\frac{1}{4n}\sum_{\begin{subarray}{c}m=-\infty\\ \end{subarray}}^{\infty}\frac{\left|q_{2m+1}\right|^{2}}{(2n+2m+1)}.$
First, let us prove that
$I_{1,1}=D_{1}(n,q),\qquad I_{1,2}=D_{2}(n,q),\qquad I_{1}=D(n,q).$ (20)
It is clear that
$Q(\pi,n)=\int_{0}^{\pi}q(t)e^{i2nt}\,dt-q_{-2n}\pi=0,\qquad Q(0,n)=0$
and that the derivative of $Q(x,n)$ with respect to $x$ is
$Q^{\prime}(x,n)=q(x)e^{i2nx}\,-q_{-2n}.$
Therefore, using the integration by parts, we see that
$\int_{0}^{\pi}Q(x,n)e^{-i(2n+2m)x}dx=\frac{\pi q_{2m}}{i(2n+2m)},$
for $n+m\neq 0$. Therefore, the Fourier decomposition of $Q(x,n)$ by the orthonormal basis $\{e^{i2mx}/\sqrt{\pi}:m\in\mathbb{Z\}}$ of $L_{2}[0,\pi]$ has the form
$Q(x,n)=Q_{n,0}+\sum_{m\neq 0,-n}\frac{q_{2m}e^{i(2n+2m)x}}{i(2n+2m)}.$
Using this decomposition in the integral for $D_{1}(n,q)$, we obtain the proof of the first equality of (20).
Now, let us prove the second equality of (20). It is clear that
$G(\pi,n)=\int_{0}^{\pi}\bigl(q(t)e^{i2nt}\,+\frac{i\pi}{2}e^{it}\,q_{-2n}\bigr)dt=\pi q_{-2n}+\frac{\pi}{2}(e^{i\pi}-1)q_{-2n}=0,$
$G(0,n)=0$ and
$G^{\prime}(x,n)=q(x)e^{i2nx}\,+\frac{i\pi}{2}e^{ix}q_{-2n}.$
Therefore,
$\int_{0}^{\pi}G(x,n)e^{-i(2n+2m+1)x}dx=\frac{\pi q_{2m+1}}{i(2n+2m+1)}.$
Thus, the Fourier decomposition of $G(x,n)$ by the orthonormal basis $\{e^{i(2m+1)x}/\sqrt{\pi}:m\in\mathbb{Z}\}$ of $L_{2}[0,\pi]$ has the form
$G(x,n)=\sum_{m\in\mathbb{Z}}\frac{q_{2m+1}e^{i(2n+2m+1)x}\,}{i(2n+2m+1)}.$
Using this decomposition of $G(x,n)$ in the integral for $D_{2}(n,q)$, we obtain the proof of the second equality of (20). The third equality of (20) follows from (19).
Now, we consider $I_{2}.$ Using the definition of $p(x)$ and the substitution t=xt=-x, we obtain
$p_{k}:=\frac{1}{\pi}\int_{0}^{\pi}p(x)e^{-ikx}dx\,=\frac{1}{\pi}\int_{0}^{\pi}q(-x)e^{-ikx}dx\,=\frac{1}{\pi}\int_{0}^{\pi}q(t)e^{ikt}dt=q_{-k}.$
Therefore, instead of $q(x)$, using $p(x)$ and repeating the proof of (20), we obtain that
$I_{2}=D(n,p),\qquad I_{2}+I_{3}=2\operatorname{Re}D(n,p).$ (21)
Thus the proof of (15) follows from (16)-(21). The lemma is proved.
Now let us consider $B_{1,n}(n^{2})$.
Lemma 2
If $q\in L_{1}[0,\pi]$, then the following formula holds
$B_{1,n}(n^{2})=-\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx,$ (22)
where $Q(x)=$ ${\displaystyle\int_{0}^{x}}q(x)dx$.
Proof. To prove (22), we estimate the left and write side of this equality separately. First estimate the right-side. Let $Q_{k}={\displaystyle\int_{0}^{\pi}}Q(x)\sin kxdx$. Then, $Q_{k}=\pi C_{k}/k$. Therefore, we have the decomposition
$Q(x)=\sum_{k=1}^{\infty}\dfrac{2}{\pi}Q_{k}\sin kx$
of $Q(x)$. Using the equality $(Q^{2}(x))^{\prime}=2q(x)Q(x)$, the decomposition of $Q(x)$, and integration by parts, and then the equality $\sin kx\sin 2nx=\frac{1}{2}(\cos(2n-k)x-\cos(2n+k)x)$, we obtain
$\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx=-\frac{1}{n}\int_{0}^{\pi}q(x)Q(x)\sin 2nxdx$
$\qquad=-\frac{2}{\pi n}\int_{0}^{\pi}q(x)\sum_{k=1}^{\infty}Q_{k}\sin kx\sin 2nxdx$
$\qquad=-\frac{1}{\pi n}\int_{0}^{\pi}q(x)\sum_{k=1}^{\infty}Q_{k}(\cos(2n-k)x-\cos(2n+k)x)dx$
$\qquad=\frac{-1}{n}\sum_{k=1}^{\infty}Q_{k}(C_{2n-k}-C_{2n+k})=\frac{-1}{n}\sum_{k=1}^{\infty}Q_{k}C_{2n-k}+\frac{1}{n}\sum_{k=1}^{\infty}Q_{k}C_{2n+k}).$
Now, using the equalities $Q_{k}=\pi C_{k}/k$, $C_{-k}=C_{k}$, and doing the substitution $k\rightarrow-k$ in the second summation of the last equality, we obtain
$\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx=\frac{-\pi}{n}\sum_{k\in\mathbb{Z},k\neq 0}\frac{C_{k}C_{2n-k}}{k}$
Thus, the right side of (22) has the form
$\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx=\frac{1}{n}\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k},$
since $C_{0}=0$.
Now, we prove that the left side of (22) also can be written in this form. Doing the substitution $k\rightarrow-k$, and then using the equality
$\frac{1}{k(2n-k)}=\frac{1}{2n}\biggl(\frac{1}{k}+\frac{1}{2n-k}\biggr),$
we see that
$B_{1,n}(n^{2})=\sum_{k\in\mathbb{Z},k\neq 0,-2n}\frac{C_{k}C_{2n+k}}{n^{2}-(n+k)^{2}}=\sum_{k\in\mathbb{Z},k\neq 0,-2n}\frac{C_{k}C_{2n+k}}{(2n+k)(-k)}=\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k(2n-k)}$
$\qquad=\frac{1}{2n}\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k}+\frac{1}{2n}\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{2n-k}.$
On the other hand, doing the substitution $k\rightarrow 2n-k$ in the second summation of the last equality, we obtain
$\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{2n-k}=\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k}.$
Therefore, we have
$B_{1,n}(n^{2})=\frac{1}{n}\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k},$
which means that (22) holds. The lemma is proved.
Now, we consider $a_{2}(n^{2})$. It is clear that $a_{2}(n^{2})=a_{2,1}(n^{2})-a_{2,2}(n^{2})$, where
$a_{2,1}(n^{2})=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{k+l}}{[n^{2}-(n+k)^{2}][n^{2}-((n+k+l))^{2}]}$
$\qquad=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(2n+k)(k+l)(2n+k+l)}$
and
$a_{2,2}(n^{2})=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(2n+k)(k+l)(2n+k+l)}.$
Using the equality
$\frac{1}{(k+l)(2n+k+l)}=\frac{1}{2n}\left(\frac{1}{k+l}+\frac{1}{2n+k+l}\right),$
for $l=0$ and $l\neq 0$, we obtain
$a_{2,1}(n^{2})=\frac{1}{4n^{2}}(S_{1}(n,q)+S_{2}(n,q)+S_{3}(n,q)+S_{4}(n,q)),$ (23)
where
$S_{1}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(k+l)},$
$S_{2}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{(2n+k)(k+l)},$
$S_{3}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(2n+k+l)},$
and
$S_{4}(n,q)=\sum\limits_{k,k+l\neq 0,2n}\frac{C_{k}C_{l}C_{-k-l}}{(2n+k)(2n+k+l)}.$
First, let us consider $S_{1}(n,q)$. Grouping the terms
$\frac{C_{k}C_{l}C_{-k-l}}{k(k+l)},\quad\frac{C_{l}C_{k}C_{-k-l}}{l(k+l)},\quad\frac{C_{k}C_{-k-l}C_{l}}{k(-l)},$
$\frac{C_{l}C_{-k-l}C_{k}}{l(-k)},\quad\frac{C_{-k-l}C_{l}C_{k}}{(-k-l)(-k)},\quad\frac{C_{-k-l}C_{k}C_{l}}{(-k-l)(-l)}$
with the equal multiplicands,
$C_{k}C_{l}C_{-k-l},\quad C_{l}C_{k}C_{-k-l},\quad C_{k}C_{-k-l}C_{l},$
$C_{l}C_{-k-l}C_{k},\quad C_{-k-l}C_{l}C_{k},\quad C_{-k-l}C_{k}q_{l}$
and using the equality
$\frac{1}{k(k+l)}+\frac{1}{l(k+l)}=\frac{1}{kl},$
we obtain that
$\sum\limits_{k,l}\frac{C_{k}C_{l}C_{-k-l}}{k(k+l)}=0,$ (24)
where the summations are taken under conditions $k,l,k+l\neq 0,$ since $C_{0}=0$. However, in summation for $S_{1}(n,q)$ it is assumed additionally that $k,k+l\neq-2n$. Therefore, we also need to consider the following expressions:
$S_{1,1}(n,q):=\frac{C_{2n}}{2n}\sum\limits_{l\neq 0,-2n}\frac{C_{2n}C_{l}C_{l-2n}}{2n(l-2n)},\qquad S_{1,2}(n,q)=\frac{C_{2n}}{2n}\sum\limits_{l\neq 0,-2n}\frac{C_{2n+l}C_{l}}{(2n+l)}.$
Using the Schwarz inequality for $l_{2}$ and taking into account that $C_{l}=O(1)$, we obtain
$S_{1,1}(n,q):=\frac{C_{2n}}{2n}O(1)=o\bigl(\frac{1}{n}\bigr),\qquad S_{1,2}(q):=\frac{C_{2n}}{2n}O(1)=o\bigl(\frac{1}{n}\bigr),$ (25)
for $q\in L_{2}[0,1]$.
Denote by $E$ the set of all functions satisfying
$C_{n}=\int_{0}^{\pi}q(x)\cos nxdx=O\bigl(\frac{1}{n}\bigr)$ (26)
as $n\rightarrow\infty$. For example, if $q$ is a function of bounded variation on $0,\pi]$, then (26) holds. There is a larger set of functions satisfying (26) (see [14]). Here we do not discuss the set $E$, since this is not the aim of this paper. We will apply the obtained further result for potential (9) which is a bounded variation.
Using (26) in (25), we obtain
$S_{1,1}(n,q)=O\bigl(\frac{1}{n^{2}}\bigr),\text{ }S_{1,2}(q)=O\bigl(\frac{1}{n^{2}}\bigr).$
These equalities with (24) imply the following:
Proposition 1
If $q\in E$, then $S_{1}(n,q)=O(n^{-2})$.
Now, we estimate $S_{j}(n,q)$, for $j=2,3,4$, in the following lemma. For this, we use the following easily checkable relations
$\sum\limits_{k\neq 0,-2n}\frac{1}{k^{2}(2n+k)}=O\bigl(\frac{1}{n}\bigr),$ (27)
$\sum\limits_{l\neq 0,-k}\frac{1}{l(k+l)^{2}}=O\bigl(\frac{1}{k}\bigr),$ (28)
and
$\sum\limits_{k\neq 0,-2n}\frac{1}{\left|k(2n+k)\right|}=O\bigl(\frac{\ln n}{n}\bigr).$ (29)
To prove (27), split the sum into three parts
$\sum\limits_{k\leq n,\text{ }k\neq 0}\frac{1}{k^{2}(2n+k)},\qquad\sum\limits_{k>n}\frac{1}{k^{2}(2n+k)},\qquad\sum\limits_{k<-n,k\neq-2n}\frac{1}{k^{2}(2n+k)}$
and estimate each of them in a standard way. Equalities (28) and (29) can be proved in the same way.
Lemma 3
If $q\in E$, then the equalities
$S_{j}(n,q)=O\bigl(\frac{1}{n}\bigr),$ (30)
for $j=2,3,4$, hold.
Proof. It follows from (26) and the definition of $S_{2}(n,q)$ that
$\left|-S_{2}(n,q)\right|=O(1)\biggl(\sum\limits_{k\neq 0,-2n}\sum\limits_{l\neq 0,-k}\frac{1}{\left|k(2n+k)l(k+l)^{2}\right|}\biggr).$
Now using (28) and then (27), we obtain
$\left|-S_{2}(n,q)\right|=O(1)\sum\limits_{k\neq 0,-2n}\frac{1}{\left|k^{2}(2n+k)\right|}=O\bigl(\frac{1}{n}\bigr).$
Now, let us prove (30) for $j=3$. By (26), we have
$\left|-S_{3}(n,q)\right|=O(1)\sum\limits_{k\neq 0,-2n}\sum\limits_{l\neq 0,-k}\frac{1}{\left|l(2n+k+l)k^{2}(k+l)\right|}.$
Making the substitution $s=k+l$, we obtain
$\left|-S_{3}(n,q)\right|=O(1)\sum\limits_{s\neq 0,-2n}\sum\limits_{k\neq 0,s}\frac{1}{\left|s(2n+s)k^{2}(s-k)\right|}.$
Now using (27) and (28), we obtain
$\left|-S_{3}(n,q)\right|=O(1)\sum\limits_{s\neq 0,2n}\frac{1}{\left|s^{2}(2n+s)\right|}=O\bigl(\frac{1}{n}\bigr),$
that is, the proof of (30) for $j=3$. Arguing as above, we conclude that
$S_{4}(n,q)=O(1)\biggl(\sum\limits_{k\neq 0,-2n}\frac{1}{\left|k(2n+k)\right|}\biggr)\biggl(\sum\limits_{s\neq 0,-2n}\frac{1}{\left|s(2n+s)\right|}\biggr).$
Therefore, (29) implies that
$S_{4}(n,q)=O\bigl(\frac{\left(\ln n\right)^{2}}{n^{2}}\bigr)=O\bigl(\frac{1}{n}\bigr).$
The lemma is proved.
It follows from Proposition 1 and Lemma 3 that
$a_{2,1}(n^{2})=O\bigl(\frac{1}{n^{3}}\bigr),$ (31)
if $q\in E$. Now let us estimate $a_{2,2}(n^{2})$. Arguing as in the proof of (23), we see that
$a_{2,2}(n^{2})=\frac{1}{4n^{2}}(H_{1}+H_{2}+H_{3}+H_{4}),$
where
$H_{1}(q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(k+l)},$
$H_{2}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{(2n+k)(k+l)},$
$H_{3}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(2n+k+l)},$
and
$H_{4}(n,q)=\sum\limits_{k,k+l\neq 0,2n}\frac{C_{k}C_{l}C_{2n+k+l}}{(2n+k)(2n+k+l)}.$
Therefore, arguing as in the proof of Lemma 3, we conclude that
$a_{2,2}(n^{2},q)=O\bigl(\frac{1}{n^{3}}\bigr),$ (32)
if $q\in E$.
Now, using Lemmas 1 and 2, (31) and (32), we obtain the following main results of this section:
Theorem 1
If $q\in E$, then the following asymptotic formula holds:
$\lambda_{n}=n^{2}-C_{2n}-\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)+O\bigl(\frac{1}{n^{3}}\bigr),$ (33)
where $D(n,f)$ and $Q(x)$ are defined in Lemma 1 and Lemma 2.
Proof. It follows from (8), (26) and (27) that
$\lambda_{n}=n^{2}+O\bigl(\frac{1}{n}\bigr).$
Using this, we easily get the estimate
$\sum_{\begin{subarray}{c}k\neq 0,-2n\end{subarray}}^{\infty}\mid\frac{1}{\lambda_{n}-(n+k)^{2}}-\frac{1}{n^{2}-(n+k)^{2}}\mid=O\bigl(\frac{1}{n^{3}}\bigr).$
By virtue of this estimate
$a_{j}\left(\lambda_{n}\right)=a_{j}\left(n^{2}\right)+O\bigl(\frac{1}{n^{3}}\bigr),$
for $j=1,2,\ldots$. Therefore it follows from (2)-(6) that
$\lambda_{n}=n^{2}-C_{2n}+a_{1}\left(n^{2}\right)+a_{2}\left(n^{2}\right)+a_{3}\left(n^{2}\right)+O\bigl(\frac{1}{n^{3}}\bigr),$ (34)
where
$a_{1}\left(n^{2}\right)=D(n,q)+2\operatorname{Re}D(n,p)+O\bigl(\frac{1}{n^{3}}\bigr)$ (35)
(see Lemma 1, Lemma 2 and (26)),
$a_{2}\left(n^{2}\right)=O\bigl(\frac{1}{n^{3}}\bigr)$ (36)
(see (31) and (32)). On the other hand, we have
$a_{3}\left(n^{2}\right)=\sum_{n_{1},n_{2},...,n_{k}=-\infty}^{\infty}\frac{C_{k}C_{l}C_{s}(C_{k+l+s}-C_{2n+k+l+s})}{k(2n+k)(k+l)(2n+k+l)(k+l+s)(2n+k+l+s)}.$
Moreover, arguing as in the proof of (36) we see that $a_{3}\left(n^{2}\right)=O\bigl(\dfrac{1}{n^{3}}\bigr)$. Thus, the proof of the theorem follows from (34)-(36).
Now, let us demonstrate the obtained results for potential (9). In this case, by direct calculation, we have
$q_{k}=\dfrac{1}{\pi}\int_{0}^{\pi}q(x)e^{-ikx}dx\,=\dfrac{1}{i\pi k}((b-a)e^{-ikc}+a-(-1)^{k}b).$
Using the equalities $Q(\pi,n)=G(\pi,n)=0$ (see the proof of Lemma 1), we obtain
$Q(x,n)=\begin{cases}\dfrac{a}{i2n}(e^{i2nx}-1)-q_{-2n}x&\text{if }x\in[0,c]\\[5.69054pt] \dfrac{b}{i2n}(e^{i2nx}-1)-q_{-2n}x+\pi q_{-2n}&\text{if }x\in(c,\pi]\end{cases}$
and
$G(x,n)=\begin{cases}\dfrac{a}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)&\text{if }x\in[0,c]\\[5.69054pt] \dfrac{b}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)+\pi q_{-2n}&\text{if }x\in(c,\pi],\end{cases}$
where $Q(x,n)$, and $G(x,n)$ are defined by (13) and (14). Therefore, by Lemma 1 we have
$D_{1}(n,q)=\frac{i}{4\pi n}\int_{0}^{\pi}q(x)Q(x,n)e^{-i2nx}\,dx-\frac{iQ_{n,0}}{4\pi n}\int_{0}^{\pi}q(x)e^{-i2nx}\,dx$
$\qquad=\frac{i}{4\pi n}\biggl(\int_{0}^{c}q(x)Q(x,n)e^{-i2nx}\,dx+\int_{c}^{\pi}q(x)Q(x,n)e^{-i2nx}\,dx\biggr)-\frac{iQ_{n,0}q_{2n}}{4n},$
and
$D_{2}(n,q)=\frac{i}{4\pi n}\int_{0}^{\pi}q(x)G(x,n)e^{-i2nx}\,dx$
$\qquad=\frac{i}{4\pi n}\biggl(\int_{0}^{c}q(x)G(x,n)e^{-i2nx}\,dx+\int_{c}^{\pi}q(x)G(x,n)e^{-i2nx}\,dx\biggr).$
We first calculate the terms in $D_{1}(n,q)$. Using $(b-a)c=b\pi$ or $ac=b(c-\pi)$ by (10) and $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$, we have
$\pi Q_{n,0}$ $=\int_{0}^{c}\bigl(\dfrac{a}{i2n}(e^{i2nx}-1)-q_{-2n}x\bigr)dx+\int_{c}^{\pi}\bigl(\dfrac{b}{i2n}(e^{i2nx}-1)-q_{-2n}x+\pi q_{-2n}\bigr)dx$
$=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}+\pi(\frac{\pi}{2}-c)q_{-2n}=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}+\pi(\frac{\pi}{2}-c)\frac{i(b-a)(e^{i2nc}-1)}{2\pi n}$
$=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}-\frac{i\pi(b+a)(e^{i2nc}-1)}{4n},$ (37)
$\frac{i}{4n}Q_{n,0}q_{2n}=\frac{i(b^{2}-a^{2})(1-\cos 2nc)}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{16\pi^{2}n^{4}},$
$\int_{0}^{c}q(x)Q(x,n)e^{-i2nx}\,dx=\int_{0}^{c}a\biggl(\dfrac{a}{i2n}(e^{i2nx}-1)-q_{-2n}x\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{a^{2}}{i2n}\int_{0}^{c}(1-e^{-i2nx})\,dx-aq_{-2n}\int_{0}^{c}xe^{-i2nx}\,dx$
$\qquad=\dfrac{a^{2}}{i2n}\biggl(c+\dfrac{e^{-i2nc}-1}{i2n}\biggr)+\dfrac{aq_{-2n}}{i2n}\biggl(ce^{-i2nc}+\dfrac{e^{-i2nc}-1}{i2n}\biggr),$
and
$\int_{c}^{\pi}q(x)Q(x,n)e^{-i2nx}\,dx=\int_{c}^{\pi}b\biggl(\dfrac{b}{i2n}(e^{i2nx}-1)-q_{-2n}x+\pi q_{-2n}\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{b^{2}}{i2n}\int_{c}^{\pi}(1-e^{-i2nx})\,dx-bq_{-2n}\int_{c}^{\pi}xe^{-i2nx}\,dx+b\pi q_{-2n}\int_{c}^{\pi}e^{-i2nx}\,dx$
$\qquad=\dfrac{b^{2}}{i2n}\biggl(\pi-c-\dfrac{e^{-i2nc}-1}{i2n}\biggr)+\dfrac{bq_{-2n}}{i2n}\biggl(\pi-ce^{-i2nc}-\dfrac{e^{-i2nc}-1}{i2n}\biggr)+\dfrac{b\pi q_{-2n}}{i2n}(e^{-i2nc}-1).$
Therefore,
$D_{1}(n,q)=\frac{i}{4\pi n}\int_{0}^{\pi}q(x)Q(x,n)e^{-i2nx}\,dx-\frac{iQ_{n,0}}{4\pi n}\int_{0}^{\pi}q(x)e^{-i2nx}\,dx$
$\qquad=\frac{i}{4\pi n}\biggl(\frac{(b^{2}-a^{2})(e^{-i2nc}-1)}{4n^{2}}-\frac{ab\pi}{i2n}+\frac{(b-a)(e^{-i2nc}-1)q_{-2n}}{4n^{2}}\biggr)-\frac{iQ_{n,0}q_{2n}}{4n}$
and substituting $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$ in the last expression, we have
$D_{1}(n,q)=-\frac{ab}{8n^{2}}+\frac{i(b^{2}-a^{2})(e^{-i2nc}-1)}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{16\pi^{2}n^{4}}$
$\qquad+\frac{i(b^{2}-a^{2})(1-\cos 2nc)}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{16\pi^{2}n^{4}}$
$\qquad=-\frac{ab}{8n^{2}}+\frac{(b^{2}-a^{2})\sin 2nc}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{8\pi^{2}n^{4}}.$ (38)
Now let us calculate the terms in $D_{2}(n,q)$:
$\int_{0}^{c}q(x)G(x,n)e^{-i2nx}\,dx=\int_{0}^{c}a\biggl(\dfrac{a}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{a^{2}}{i2n}\int_{0}^{c}(1-e^{-i2nx})\,dx+a\dfrac{\pi}{2}q_{-2n}\int_{0}^{c}xe^{-i(2n-1)x}\,dx-a\dfrac{\pi}{2}q_{-2n}\int_{0}^{c}xe^{-i2nx}\,dx$
$\qquad=\dfrac{a^{2}}{i2n}\biggl(c+\dfrac{e^{-i2nc}-1}{i2n}\biggr)-\dfrac{a\pi q_{-2n}}{i2(2n-1)}\biggl(ce^{-i(2n-1)c}+\dfrac{e^{-i(2n-1)c}-1}{i(2n-1)}\biggr)$
$\qquad+\dfrac{a\pi q_{-2n}}{i4n}\biggl(ce^{-i2nc}+\dfrac{e^{-i2nc}-1}{i2n}\biggr)$
and
$\int_{c}^{\pi}q(x)G(x,n)e^{-i2nx}\,dx=\int_{c}^{\pi}b\biggl(\dfrac{b}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)+\pi q_{-2n}\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{b^{2}}{i2n}\int_{c}^{\pi}(1-e^{-i2nx})\,dx+b\dfrac{\pi}{2}q_{-2n}\int_{c}^{\pi}xe^{-i(2n-1)x}\,dx-b\dfrac{\pi}{2}q_{-2n}\int_{c}^{\pi}xe^{-i2nx}\,dx+b\pi q_{-2n}\int_{c}^{\pi}e^{-i2nx}\,dx$
$\qquad=\dfrac{b^{2}}{i2n}\biggl(\pi-c-\dfrac{e^{-i2nc}-1}{i2n}\biggr)+\dfrac{b\pi q_{-2n}}{i2(2n-1)}\biggl(\pi+ce^{-i(2n-1)c}+\dfrac{e^{-i(2n-1)c}+1}{i(2n-1)}\biggr)$
$\qquad+\dfrac{b\pi q_{-2n}}{i4n}\biggl(\pi-ce^{-i2nc}-\dfrac{e^{-i2nc}-1}{i2n}\biggr)+\dfrac{b\pi q_{-2n}}{i2n}(e^{-i2nc}-1)$
give
$D_{2}(n,q)=\frac{i}{4\pi n}\int_{0}^{\pi}q(x)G(x,n)e^{-i2nx}\,dx$
$\qquad=\frac{i}{4\pi n}\biggl(\frac{(b^{2}-a^{2})(e^{-i2nc}-1)}{4n^{2}}-\frac{ab\pi}{i2n}+\frac{b\pi^{2}(1+e^{-i(2n-1)c})q_{-2n}}{i2(2n-1)}+\frac{b\pi^{2}(1-e^{-i2nc})q_{-2n}}{i4n}$
$\qquad-\frac{\pi(b-a)(1-e^{-i2nc})q_{-2n}}{8n^{2}}-\frac{\pi q_{-2n}(b+a+(b-a)e^{-i(2n-1)c})}{2(2n-1)^{2}}\biggr).$
Using $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$, we obtain
$D_{2}(n,q)=-\frac{ab}{8n^{2}}+\frac{i(b^{2}-a^{2})(e^{-i2nc}-1)}{16\pi n^{3}}-\frac{ib(b-a)(1-e^{-i2nc})(e^{i2nc}+e^{ic})}{16n^{2}(2n-1)}+\frac{ib(b-a)(\cos 2nc-1)}{16n^{3}}$
$\qquad+\frac{(b-a)^{2}(\cos 2nc-1)}{32\pi n^{4}}+\frac{(b-a)(e^{-i2nc}-1)[(b+a)e^{i2nc}+(b-a)e^{ic}]}{16\pi n^{2}(2n-1)^{2}}.$ (39)
Thus, by (38), (39), and Lemma 1, we write
$D(n,q)=D_{1}(n,q)+D_{2}(n,q)=-\frac{ab}{4n^{2}}+\frac{(b^{2}-a^{2})\sin 2nc}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{8\pi^{2}n^{4}}$
$\qquad+\frac{i(b^{2}-a^{2})(e^{-i2nc}-1)}{16\pi n^{3}}+\frac{ib(b-a)(\cos 2nc-1)}{16n^{3}}-\frac{ib(b-a)(1-e^{-i2nc})(e^{i2nc}+e^{ic})}{16n^{2}(2n-1)}$
$\qquad+\frac{(b-a)^{2}(\cos 2nc-1)}{32\pi n^{4}}+\frac{(b-a)(e^{-i2nc}-1)[(b+a)e^{i2nc}+(b-a)e^{ic}]}{16\pi n^{2}(2n-1)^{2}}$
or
$D(n,q)=D_{1}(n,q)+D_{2}(n,q)=-\frac{ab}{4n^{2}}+\frac{(b^{2}-a^{2})\sin 2nc}{8\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{8\pi^{2}n^{4}}$
$\qquad+\frac{i(b^{2}-a^{2})(\cos 2nc-1)}{16\pi n^{3}}+\frac{ib(b-a)(\cos 2nc-1)}{16n^{3}}-\frac{ib(b-a)(1-e^{-i2nc})(e^{i2nc}+e^{ic})}{16n^{2}(2n-1)}$
$\qquad+\frac{(b-a)^{2}(\cos 2nc-1)}{32\pi n^{4}}+\frac{(b-a)(e^{-i2nc}-1)[(b+a)e^{i2nc}+(b-a)e^{ic}]}{16\pi n^{2}(2n-1)^{2}}.$ (40)
Now, let us calculate $\operatorname{Re}D(n,p)$, where $p(x)$ is defined in Lemma 1. For the Kronig-Penney potential, $p(x)$ becomes
$p(x)=\begin{cases}b&\text{if }x\in[0,\pi-c),\\[5.69054pt] a&\text{if }x\in[\pi-c,\pi]\end{cases}.$
Therefore,
$D_{1}(n,p)=\frac{i}{4\pi n}\int_{0}^{\pi}p(x)Q(x,n)e^{-i2nx}\,dx-\frac{iQ_{n,0}}{4\pi n}\int_{0}^{\pi}p(x)e^{-i2nx}\,dx$
$\qquad=\frac{i}{4\pi n}\biggl(\int_{0}^{\pi-c}p(x)Q(x,n)e^{-i2nx}\,dx+\int_{\pi-c}^{\pi}p(x)Q(x,n)e^{-i2nx}\,dx\biggr)-\frac{iQ_{n,0}p_{2n}}{4n}$
and
$D_{2}(n,p)=\frac{i}{4\pi n}\int_{0}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx$
$\qquad=\frac{i}{4\pi n}\biggl(\int_{0}^{\pi-c}p(x)G(x,n)e^{-i2nx}\,dx+\int_{\pi-c}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx\biggr).$
We first calculate the terms in $D_{1}(n,p)$. Using $(b-a)c=b\pi$ or $ac=b(c-\pi)$ by (10), $p_{2n}=q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$, and (37), we have
$\frac{i}{4n}Q_{n,0}p_{2n}=-\frac{i(b^{2}-a^{2})(e^{i2nc}-1)^{2}}{32\pi n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{32\pi^{2}n^{4}},$
$\int_{0}^{\pi-c}p(x)Q(x,n)e^{-i2nx}\,dx=\int_{0}^{\pi-c}b\biggl(\dfrac{a}{i2n}(e^{i2nx}-1)-q_{-2n}x\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\int_{0}^{\pi-c}(1-e^{-i2nx})\,dx-bq_{-2n}\int_{0}^{\pi-c}xe^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\biggl(\pi-c+\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{bq_{-2n}}{i2n}\biggl((\pi-c)e^{i2nc}+\dfrac{e^{i2nc}-1}{i2n}\biggr),$
and
$\int_{\pi-c}^{\pi}p(x)Q(x,n)e^{-i2nx}\,dx=\int_{\pi-c}^{\pi}a\biggl(\dfrac{b}{i2n}(e^{i2nx}-1)-q_{-2n}x+\pi q_{-2n}\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\int_{\pi-c}^{\pi}(1-e^{-i2nx})\,dx-aq_{-2n}\int_{\pi-c}^{\pi}xe^{-i2nx}\,dx+a\pi q_{-2n}\int_{\pi-c}^{\pi}e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\biggl(c-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{aq_{-2n}}{i2n}\biggl(\pi-(\pi-c)e^{i2nc}-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{a\pi q_{-2n}}{i2n}(e^{i2nc}-1).$
Therefore,
$D_{1}(n,p)=\frac{i}{4\pi n}\int_{0}^{\pi}p(x)Q(x,n)e^{-i2nx}\,dx-\frac{iQ_{n,0}}{4\pi n}\int_{0}^{\pi}p(x)e^{-i2nx}\,dx$
$\qquad=\frac{i}{4\pi n}\biggl(\frac{ab\pi}{i2n}+\frac{(a-b)(e^{i2nc}-1)q_{-2n}}{4n^{2}}-\frac{a\pi(e^{i2nc}-1)q_{-2n}}{i2n}\biggr)-\frac{iQ_{n,0}p_{2n}}{4n}$
and substituting $p_{2n}=q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$ in the last expression, we have
$D_{1}(n,p)=\frac{ab}{8n^{2}}-\frac{i(b-a)(b+3a)(e^{i2nc}-1)^{2}}{32\pi n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{16\pi^{2}n^{4}}.$ (41)
Now let us calculate the terms in $D_{2}(n,p)$:
$\int_{0}^{\pi-c}p(x)G(x,n)e^{-i2nx}\,dx=\int_{0}^{\pi-c}b\biggl(\dfrac{a}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\int_{0}^{\pi-c}(1-e^{-i2nx})\,dx+b\dfrac{\pi}{2}q_{-2n}\int_{0}^{\pi-c}xe^{-i(2n-1)x}\,dx-b\dfrac{\pi}{2}q_{-2n}\int_{0}^{\pi-c}xe^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\biggl(\pi-c+\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{b\pi q_{-2n}}{i2(2n-1)}\biggl((\pi-c)e^{i(2n-1)c}+\dfrac{e^{i(2n-1)c}+1}{i(2n-1)}\biggr)$
$\qquad+\dfrac{b\pi q_{-2n}}{i4n}\biggl((\pi-c)e^{i2nc}+\dfrac{e^{i2nc}-1}{i2n}\biggr)$
and
$\int_{\pi-c}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx=\int_{\pi-c}^{\pi}a\biggl(\dfrac{b}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)+\pi q_{-2n}\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\int_{\pi-c}^{\pi}(1-e^{-i2nx})\,dx+a\dfrac{\pi}{2}q_{-2n}\int_{\pi-c}^{\pi}xe^{-i(2n-1)x}\,dx-a\dfrac{\pi}{2}q_{-2n}\int_{\pi-c}^{\pi}xe^{-i2nx}\,dx+a\pi q_{-2n}\int_{\pi-c}^{\pi}e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\biggl(c-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{a\pi q_{-2n}}{i2(2n-1)}\biggl(\pi-(\pi-c)e^{i(2n-1)c}+\dfrac{-e^{i(2n-1)c}+1}{i(2n-1)}\biggr)$
$\qquad+\dfrac{a\pi q_{-2n}}{i4n}\biggl(\pi-(\pi-c)e^{i2nc}-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{a\pi q_{-2n}}{i2n}(e^{i2nc}-1)$
give
$D_{2}(n,p)=\frac{i}{4\pi n}\int_{0}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx$
$\qquad=\frac{i}{4\pi n}\biggl(\frac{ab\pi}{i2n}+\frac{a\pi^{2}(1-e^{i(2n-1)c})q_{-2n}}{i2(2n-1)}-\frac{\pi(b+a+(b-a)e^{i(2n-1)c})q_{-2n}}{2(2n-1)^{2}}$
$\qquad+\frac{a\pi^{2}(1-e^{i2nc})q_{-2n}}{i4n}+\frac{\pi(b-a)(1-e^{i2nc})q_{-2n}}{8n^{2}}+\dfrac{a\pi(e^{i2nc}-1)q_{-2n}}{i2n}\biggr).$
Using $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$, we obtain
$D_{2}(n,p)=\frac{ab}{8n^{2}}+\frac{i(b-a)(1-e^{i(2n-1)c})(e^{i2nc}-1)}{16n^{2}(2n-1)}+\frac{(b^{2}-a^{2}+(b-a)^{2}e^{i(2n-1)c})(e^{i2nc}-1)}{16\pi n^{2}(2n-1)^{2}}$
$\qquad-\frac{ia(b-a)(e^{i2nc}-1)^{2}}{32n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{64\pi n^{4}}+\frac{ia(b-a)(e^{i2nc}-1)^{2}}{16\pi n^{3}}.$ (42)
Thus, by (41), (42), and Lemma 1, we write
$D(n,p)=D_{1}(n,p)+D_{2}(n,p)=\frac{ab}{4n^{2}}-\frac{i(b-a)(b+3a)(e^{i2nc}-1)^{2}}{32\pi n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{16\pi^{2}n^{4}}$
$\qquad+\frac{i(b-a)(1-e^{i(2n-1)c})(e^{i2nc}-1)}{16n^{2}(2n-1)}+\frac{(b^{2}-a^{2}+(b-a)^{2}e^{i(2n-1)c})(e^{i2nc}-1)}{16\pi n^{2}(2n-1)^{2}}$
$\qquad-\frac{ia(b-a)(e^{i2nc}-1)^{2}}{32n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{64\pi n^{4}}+\frac{ia(b-a)(e^{i2nc}-1)^{2}}{16\pi n^{3}}.$
and simplifying a bit
$D(n,p)=D_{1}(n,p)+D_{2}(n,p)=\frac{ab}{4n^{2}}+\frac{i(a^{2}-b^{2})(e^{i2nc}-1)^{2}}{32\pi n^{3}}$
$\qquad+\frac{i(b-a)(1-e^{i(2n-1)c})(e^{i2nc}-1)}{16n^{2}(2n-1)}-\frac{ia(b-a)(e^{i2nc}-1)^{2}}{32n^{3}}$
$\qquad+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{64\pi n^{4}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{16\pi^{2}n^{4}}+\frac{(b^{2}-a^{2}+(b-a)^{2}e^{i(2n-1)c})(e^{i2nc}-1)}{16\pi n^{2}(2n-1)^{2}}.$ (43)
Now we calculate $B_{1,n}(n^{2})$ defined by (22) in Lemma 2 . Expressing $Q(x)$ defined in Lemma 2 for the Kronig-Penney potential as
$Q(x)=\begin{cases}ax&\text{if }x\in[0,c],\\[5.69054pt] bx-b\pi&\text{if }x\in(c,\pi]\end{cases},$
by direct calculation, we obtain
$B_{1,n}(n^{2})=-\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx=\frac{ab\cos 2nc}{2n^{2}}-\frac{(b^{2}-a^{2})\sin 2nc}{4\pi n^{3}}.$ (44)
Therefore, by
$C_{k}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos kxdx=\frac{a-b}{\pi k}\sin kc,$
(40), (43), and (44), we can calculate:
$-C_{2n}-\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)$
$\qquad=-\frac{a-b}{2\pi n}\sin 2nc+\frac{ab\cos 2nc}{2n^{2}}+\frac{ab}{4n^{2}}+O\bigl(\frac{1}{n^{3}}\bigr).$
Thus, by (33), we have the following result:
Theorem 2
If $q(x)$ has the form (9), then the following asymptotic formula holds:
$\lambda_{n}=n^{2}+\frac{b-a}{2\pi n}\sin 2nc+\frac{ab(2\cos 2nc+1)}{4n^{2}}+O\bigl(\frac{1}{n^{3}}\bigr).$
3 Estimates for the Small Eigenvalues
For simplicity of reading, we first give the main ideas of the proofs of the main results of this section. To give estimates for the small Dirichlet eigenvalues, first, we prove (See Theorem 1) that Dirichlet eigenvalues satisfy the equation
$\lambda-n^{2}+C_{2n}-\sum\limits_{k=1}^{\infty}a_{k}(\lambda)=0,$ (45)
in the set $I_{n}:=[n^{2}-M,n^{2}+M]$, where $M=\max\{|a|,b\}$, $C_{k}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos kxdx$ and the infinite series $a_{k}$ is defined by (5), if the following condition holds:
Condition 1
To consider the $n$th eigenvalue $\lambda_{n}$ for $n>1,$ we assume that the condition M<(2n1)/2M<(2n-1)/2 holds. In case $n=1$, we assume that the condition $M\leq 1/2$ holds.
Then, to use numerical methods we take finite sums instead of the infinite series in the equations obtained. To approximate the roots of the equation (45), we use the fixed point iteration. It can also be used the Newton-Raphson method but in this case it is necessary to compute the derivative of the function $K_{n}(\lambda)$ defined by (55). Then, using the Banach fixed point theorem, we prove that each of these equations containing the finite sums has a unique solution in the appropriate set $I_{n}$ (see Theorem 4). Moreover, we give error analysis (see Theorem 4 and Theorem 5) and present a numerical example.
Now, we state some preliminary facts. It is well known that the spectrum of the operator $L(q)$ is discrete and for large enough $n$, there is one eigenvalue (counting with multiplicity) in the neighborhood of $n^{2}$. See the basic and detailed classical results in [1, 2, 4, 5] and references therein. The eigenvalues of the operators $L(0)$ are $n^{2}$, for $n\in\mathbb{Z}^{+}$ and all the eigenvalues of $L(0)$ are simple.
It is also known that [6]
$|\lambda_{n}-n^{2}|\leq M,$
for $n\geq 1$. Therefore, we have
$n^{2}-M\leq\lambda_{n}\leq n^{2}+M,$ (46)
for $n\geq 1$. If $k\neq\pm n$, then
$|\lambda_{n}-k^{2}|\geq|n^{2}-k^{2}|-M=|n-k||n+k|-M$
$\qquad\geq(2n-1)-M,$ (47)
for $n\geq 1$ and under the assumptions about $M$ given in Condition 1. In particular, if $n=1$, we have $|\lambda_{1}|\leq 1+M$ and
$|\lambda_{1}-k^{2}|\geq||\lambda_{1}|-k^{2}|\geq 4-|\lambda_{1}|\geq 3-M,$
for $|k|\geq 2$. Besides, if $n\geq 2$, we have $|\lambda_{n}|\geq|\lambda_{2}|\geq 4-M$ and
$|\lambda_{n}-k^{2}|\geq||\lambda_{2}|-k^{2}|\geq|\lambda_{2}|-1\geq 3-M,$
for $k\neq\pm n$.
We stress that, the iteration formula (2) was used in [13] for large eigenvalues to obtain asymptotic formulas. In this section, we find conditions on potential (9) for which the iteration formula (2) is also valid for the small eigenvalues, as $m$ tends to infinity. We also note that, it is not easy to give such conditions, there are many technical calculations.
We remind that, since the potential $q$ is of the form (9), we have
$C_{k}=\frac{a-b}{\pi k}\sin kc,\qquad C_{-k}=C_{k},$
for $k=1,2,\ldots$. Now, in order to give the main results, we prove the following lemma. Without loss of generality, we assume that $\Psi_{n}(x)$ is a normalized eigenfunction corresponding to the eigenvalue $\lambda_{n}$.
Lemma 4
If the assumptions about $M$ given in Condition 1 hold, then the statements
(a) $\lim_{m\rightarrow\infty}R_{m}(\lambda)=0,$ and (b) $|(\Psi_{n},\sin nx)|>0$
are valid, where $R_{m}(\lambda)$ is defined by (6).
Proof. (a) Since $\Psi_{n}||=1$ and $\sin kx||=\sqrt{\pi}/\sqrt{2}$, by the Schwarz inequality, we have $|(q\Psi_{n},\sin kx)|\leq M\sqrt{\pi}/\sqrt{2}$. Considering the number of terms and the greatest summands of $R_{2m}(\lambda)$ in absolute value, by (46) and (47) we obtain,
$|R_{2m}(\lambda_{1})|<\frac{4^{m}|C_{1}||C_{2}|^{2m}M\sqrt{\pi}/\sqrt{2}}{(2n-1-M)^{2m+1}}<\frac{(b-a)}{\sqrt{2\pi}}(\frac{2}{\pi})^{2m}.$
Therefore, $\lim_{m\rightarrow\infty}R_{m}(\lambda)=0$.
(b) Suppose the contrary, $(\Psi_{n},\sin nx)=0$. Since the system of root functions
$\{\sqrt{2}\sin kx/\sqrt{\pi}:k\in\mathbb{Z}^{+}\}$
of $L(0)$ forms an orthonormal basis for $L_{2}[0,\pi]$, we have the decomposition
$\dfrac{\pi}{2}\Psi_{n}=(\Psi_{n},\sin nx)\sin nx+\sum_{k\in\mathbb{Z}^{+},k\neq n}\left(\Psi_{n},\sin kx\right)\sin kx$
for the normalized eigenfunction $\Psi_{n}$ corresponding to the eigenvalue $\lambda_{n}$ of $L(q)$. By Parseval’s equality, we obtain
$\sum_{k\in\mathbb{Z}^{+},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\frac{\pi}{2}.$
Using the relation,
$(\lambda_{N}-n^{2})(\Psi_{N},\sin nx)=(q\Psi_{N},\sin nx)$
which is obtained from
$\Psi_{N}^{\prime\prime}(x)+q(x)\Psi_{N}(x)=\lambda_{N}\Psi_{N}(x),$
by multiplying both sides of the equality by $\sin nx$, where $\Psi_{N}(x)$ is an eigenfunction corresponding to the eigenvalue $\lambda_{N}$, the Bessel inequality and (47), we obtain
$\sum_{k\in\mathbb{Z}^{+},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\sum_{k\in\mathbb{Z}^{+},k\neq n}\frac{|(q\Psi_{n},\sin kx)|^{2}}{|\lambda_{n}-k^{2}|^{2}}$
$\qquad\leq\frac{1}{(2n-1-M)^{2}}\sum_{k\in\mathbb{Z},k\neq\pm n}|(q\Psi_{n},\sin kx)|^{2}<\frac{\pi(M)^{2}}{2(2n-1-M)^{2}}<\frac{\pi}{2},$
which contradicts $\sum_{k\in\mathbb{Z}^{+},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\pi/2$ and completes the proof.
Before stating the main results, we want to approximate the following series defined in Lemma 1 in a different way.
$A_{1,n}(n^{2})$ $=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}{}^{2}}{n^{2}-(n+k)^{2}}=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}^{2}}{(2n+k)(-k)}=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq-2n\end{subarray}}^{-1}\frac{C_{k}^{2}}{(2n+k)(-k)}+\sum_{k=1}^{\infty}\frac{C_{k}^{2}}{(2n+k)(-k)}$
$=\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\biggl(\frac{C_{k}^{2}}{k(2n-k)}-\frac{C_{k}^{2}}{k(2n+k)}\biggr)=\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{2C_{k}^{2}}{(2n-k)(2n+k)}.$ (48)
To this end, we introduce the integral
$I=\int_{0}^{\pi}(P(x,n)-\frac{C_{4n}}{2n}\sin 2nx)^{2}\cos 4nxdx,$ (49)
where
$P(x,n)=\int_{0}^{x}q(t)\cos 2ntdt-C_{2n}x.$
It is obvious that
$P(0,n)=P(\pi,n)=0,\qquad P^{\prime}(x,n)=q(x)\cos 2nx-C_{2n},$
and
$P(x,n)=\begin{cases}\dfrac{a}{2n}\sin 2nx-C_{2n}x&\text{if }x\in[0,c],\\[5.69054pt] \dfrac{b}{2n}\sin 2nx-C_{2n}x+C_{2n}\pi&\text{if }x\in(c,\pi].\end{cases}$
The Fourier sine coefficients of $P(x,n)$ are
$P_{k}(n)=\int_{0}^{\pi}P(x,n)\sin kxdx=\frac{\pi(C_{2n+k}+C_{2n-k})}{2k},\qquad P_{2n}(n)=\frac{\pi C_{4n}}{4n}.$
Using the Fourier decomposition
$P(x,n)-\frac{C_{4n}}{2n}\sin 2nx=\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{C_{2n+k}+C_{2n-k}}{k}\sin kx$
of $P(x,n)-\dfrac{C_{4n}}{2n}\sin 2nx$ in the integral $I$ defined by (49), we obtain
$I$ $=\int_{0}^{\pi}(P(x,n)-\frac{C_{4n}}{2n}\sin 2nx)^{2}\cos 4nxdx=\int_{0}^{\pi}\biggl(\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{C_{2n+k}+C_{2n-k}}{k}\sin kx\biggr)^{2}\cos 4nxdx$
$=\int_{0}^{\pi}\biggl(\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{C_{4n-k}+C_{k}}{2n-k}\sin(2n-k)x\frac{C_{4n+k}+C_{k}}{2n+k}\sin(2n+k)x\biggr)\cos 4nxdx$
$=-\frac{\pi}{4}\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{(C_{4n-k}+C_{k})(C_{4n+k}+C_{k})}{(2n-k)(2n+k)}$
and therefore
$\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{(C_{4n-k}+C_{k})(C_{4n+k}+C_{k})}{(2n-k)(2n+k)}=\frac{-4I}{\pi}.$ (50)
Now, the integration by parts in the integral $I$ defined by (49) gives
$I$ $=-\frac{1}{2n}\int_{0}^{\pi}P(x,n)(q(x)\cos 2nx-C_{2n})\sin 4nxdx+\frac{C_{4n}}{4n^{2}}\int_{0}^{\pi}\sin 2nx(q(x)\cos 2nx-C_{2n})\sin 4nxdx$
$=-\frac{1}{4n}\int_{0}^{\pi}P(x,n)q(x)[\sin 6nx+\sin 2nx]dx+\frac{C_{2n}}{2n}\int_{0}^{\pi}P(x,n)\sin 4nxdx+\frac{C_{4n}}{8n^{2}}\int_{0}^{\pi}q(x)[1-\cos 8nx]dx$
$=-\frac{1}{4n}I_{1}+\frac{C_{2n}}{2n}I_{2}-\frac{\pi C_{4n}C_{8n}}{8n^{2}},$
where
$I_{1}=\int_{0}^{\pi}P(x,n)q(x)[\sin 6nx+\sin 2nx]dx,\qquad I_{2}=\int_{0}^{\pi}P(x,n)\sin 4nxdx.$
By direct calculations, we obtain
$I_{1}$ $=-\frac{ab\pi}{4n}+\frac{(b^{2}-a^{2})\sin 8nc}{32n^{2}}-\frac{(b-a)^{2}}{16\pi n^{3}}+\frac{(b-a)^{2}\cos 4nc}{18\pi n^{3}}+\frac{(b-a)^{2}\cos 8nc}{144\pi n^{3}},$
$I_{2}$ $=\frac{(b-a)\sin 2nc}{16n^{2}}-\frac{(b-a)\sin 6nc}{48n^{2}},$
and
$I$ $=\frac{ab\pi}{16n^{2}}-\frac{(b^{2}-a^{2})\sin 8nc}{128n^{3}}+\frac{13(b-a)^{2}\cos 4nc}{3.2^{9}\pi n^{4}}$
$+\frac{(b-a)^{2}}{128\pi n^{4}}-\frac{5(b-a)^{2}\cos 8nc}{9.2^{7}\pi n^{4}}+\frac{(b-a)^{2}\cos 12nc}{2^{9}\pi n^{4}}.$
Therefore, by (48)-(50) we have
$|A_{1,n}(n^{2})|$ $<\frac{12}{\pi}|I|=\frac{3ab}{4n^{2}}-\frac{3(b^{2}-a^{2})\sin 8nc}{32\pi n^{3}}$ (51)
$+\frac{13(b-a)^{2}\cos 4nc}{2^{7}\pi^{2}n^{4}}+\frac{3(b-a)^{2}}{32\pi^{2}n^{4}}-\frac{5(b-a)^{2}\cos 8nc}{96\pi^{2}n^{4}}+\frac{3(b-a)^{2}\cos 12nc}{2^{7}\pi^{2}n^{4}}.$
Now, letting $m$ tend to infinity in equation (2), we obtain the following main results.
Theorem 3
If the assumptions about $M$ given in Condition 1 hold, then $\lambda_{n}$ is an eigenvalue of $L(q)$ if and only if it is the root of the equation
$\lambda-n^{2}+C_{2n}-\sum\limits_{k=1}^{\infty}a_{k}(\lambda)=0$ (52)
in the set $I_{n}:=[n^{2}-M,n^{2}+M]$.
Proof. (a) By Lemma 4, letting $m$ tend to infinity in equation (2), we obtain
$\lambda_{n}-n^{2}+C_{2n}-\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})=0,$
where $C_{k}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos kxdx$.
Now, we prove that the root of (52) lying in the interval $I_{n}$ is an eigenvalue of the operator $L$. To do this, we first give the following relations:
$a_{1}(\lambda)=a_{1}(n^{2})+\sum_{\begin{subarray}{c}n_{1}=-\infty\\ n_{1}\neq 0,2n\end{subarray}}^{\infty}\frac{C_{n_{1}}(C_{n_{1}}-C_{2n+n_{1}})(\lambda-n^{2})}{n_{1}(2n+n_{1})[\lambda-(n+n_{1})^{2}]}$
and
$a_{1}(n^{2})=A_{1,n}(n^{2})+B_{1,n}(n^{2}),$ (53)
where $A_{1,n}(n^{2})$ and $B_{1,n}(n^{2})$ are given by (11) and (12), respectively.
The equation
$F_{n}(\lambda):=\lambda-n^{2}+C_{2n}-A_{1,n}(n^{2})-B_{1,n}(n^{2})$
has one root in the interval $I_{n}=[n^{2}-M,n^{2}+M]$ and
$|F_{n}(\lambda)|\geq|\lambda-n^{2}|-|C_{2n}|-|A_{1,n}(n^{2})|-|B_{1,n}(n^{2})|.$
Estimating $|A_{1,n}(n^{2})|$ and $|B_{1,n}(n^{2})|$ by (51) and (44), we have
$|F_{n}(\lambda)|>0.16,$
for all $\lambda$ from the boundary of $I_{n}$, for $n=1$. Now we define
$G_{n}(\lambda):=\lambda-n^{2}+C_{2n}-\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n}).$
Estimating the summands of $|a_{1}(\lambda)-a_{1}(n^{2})|$ and $|a_{k}(\lambda)|$ by considering the greatest summands of them, for $n=1$, we obtain
$|a_{1}(\lambda)-a_{1}(n^{2})|<\frac{16}{45\pi^{2}},\qquad|a_{k}(\lambda)|<\frac{\sqrt{2}^{k}}{\pi^{k+1}},$
for $k\geq 2$, where $a_{1}(n^{2})$ is defined by (53). Therefore, it follows by the geometric series formula that
$|a_{1}(\lambda)-a_{1}(n^{2})|+\sum\limits_{k=2}^{\infty}|a_{k}(\lambda)|<\frac{16}{45\pi^{2}}+\frac{\pi+1}{\pi^{2}(\pi-1)}<0.1534$
for $n=1$. Hence
$|G_{n}(\lambda)-F_{n}(\lambda)|=\biggl|a_{1}(\lambda)-a_{1}(n^{2})+\sum\limits_{k=2}^{\infty}a_{k}(\lambda)\biggr|$
$\qquad\leq|a_{1}(\lambda)-a_{1}(n^{2})|+\sum\limits_{k=2}^{\infty}|a_{k}(\lambda)|$
$\qquad<\frac{16}{45\pi^{2}}+\frac{\pi+1}{\pi^{2}(\pi-1)}<0.1534$
for all $\lambda$ from the boundary of $I_{n}$, for $n=1$. Similar estimations can be obtained also for $n\geq 2$. Therefore $|G_{n}(\lambda)-F_{n}(\lambda)|<|F_{n}(\lambda)|$ holds for all $\lambda$ from the boundary of $I_{n}$ and by Rouche’s theorem, $G_{n}(\lambda)$ has one root in the set $I_{n}$, for $n\geq 1$. Hence, $L(q)$ has one eigenvalue (counting multiplicity) in $I_{n}$, which is the root of (52). On the other hand, equation (52) has exactly one root (counting multiplicity) in $I_{n}$. Thus, $\lambda\in I_{n}$ is an eigenvalue of $L(q)$ if and only if, it is the root of (52).
Now, we can use numerical methods by taking finite sums instead of the infinite series in (52) and obtain
$\lambda-n^{2}+C_{2n}-\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda)=0,$
where
$a_{r,k,n}(\lambda)=\sum_{n_{1},n_{2},...,n_{k}=-r}^{r}\frac{C_{n_{1}}C_{n_{2}}\cdots C_{n_{k}}(C_{n_{1}+n_{2}+\cdots+n_{k}}-C_{n_{1}+n_{2}+\cdots+n_{k}+2n})}{[\lambda-(n+n_{1})^{2}]\cdots[\lambda-(n+n_{1}+\cdots+n_{k})^{2}]}.$ (54)
Define the function
$K_{n}(\lambda):=\lambda-n^{2}-g_{n}(\lambda),$ (55)
where
$g_{n}(\lambda)=-C_{2n}+\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda).$ (56)
Then,
$\lambda=n^{2}+g_{n}(\lambda),$ (57)
for $n\geq 1$.
Now we state another main result.
Theorem 4
Suppose that the assumptions about $M$ given in Condition 1 hold. Then for all $x$ and $y$ from the interval $I_{n}=[n^{2}-M,n^{2}+M]$, the relations
$|g_{n}(x)-g_{n}(y)|\leq L_{n}|x-y|,$ (58)
$L_{n}=\frac{9(b-a)^{2}}{4\pi(2n-1-M)[4\pi(2n-1-M)-3(b-a)]}\leq\frac{9}{2\pi(2\pi-3)}<1,$
hold and equation (57) has a unique solution $\rho_{n}$ in $I_{n}$. Moreover
$|\lambda_{n}-\rho_{n}|<\frac{(b-a)^{s+2}}{2M\sqrt{2}^{s}\pi^{s+1}(2n-1-M)^{s}[\sqrt{2}\pi(2n-1-M)-(b-a)](1-L_{n})}$
$\qquad+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M](1-L_{n})}.$ (59)
Proof. First we prove (58) by using the mean-value theorem. To do this, we estimate $|g_{n}^{\prime}(\lambda)|=|\dfrac{d}{d\lambda}g_{n}(\lambda)|$. By (56), we have
$|g_{n}^{\prime}(\lambda)|=\biggl|\sum\limits_{k=1}^{s}\frac{d}{d\lambda}a_{r,k,n}(\lambda)\biggr|\leq\sum\limits_{k=1}^{s}\bigl|\frac{d}{d\lambda}a_{r,k,n}(\lambda)\bigr|.$
First let us estimate the summands of the term $|\dfrac{d}{d\lambda}a_{r,k,n}(\lambda)|$:
$\biggl|\frac{d}{d\lambda}(a_{r,k,n}(\lambda))\biggr|=\biggl|\sum_{n_{1},n_{2},...,n_{k}=-r}^{r}\frac{d}{d\lambda}\frac{C_{n_{1}}C_{n_{2}}\cdots C_{n_{k}}(C_{n_{1}+n_{2}+\cdots+n_{k}}-C_{n_{1}+n_{2}+\cdots+n_{k}+2n})}{[\lambda-(n+n_{1})^{2}]\cdots[\lambda-(n+n_{1}+\cdots+n_{k})^{2}]}\biggr|$
$\qquad<\frac{3^{k+1}(b-a)^{k+1}}{4^{k+1}\pi^{k+1}(2n-1-M)^{k+1}}\leq\bigl|\frac{3}{2\pi}\bigr|^{k+1},$
for $k\geq 1$. Thus, by the geometric series formula, we obtain
$\sum\limits_{k=1}^{r}\bigl|\frac{d}{d\lambda}a_{r,k,n}(\lambda)\bigr|<\frac{9}{2\pi(2\pi-3)}.$
Hence,
$|g_{n}^{\prime}(\lambda)|<\frac{9(b-a)^{2}}{4\pi(2n-1-M)[4\pi(2n-1-M)-3(b-a)]}=L_{n}\leq\frac{9}{2\pi(2\pi-3)}<1.$ (60)
Since the inequality
$|g_{n}^{\prime}(\lambda)|<L_{n}<1$ (61)
holds for all $x,y\in I_{n}$, (58) holds by the mean value theorem and equation (57) has a unique solution $\rho_{n}$ in $I_{n}$, by the contraction mapping theorem. Now let us prove (59). By (55), we have $K_{n}(x)=x-n^{2}-g_{n}(x)$ and by the definition of $\rho_{n}$, we write $K_{n}(\rho_{n})=0$. Therefore by (52) and (57), we obtain
$|K_{n}(\lambda_{n})-K_{n}(\rho_{n})|=|K_{n}(\lambda_{n})|$
$\qquad=\biggl|\lambda_{n}-n^{2}+C_{2n}-\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|$
$\qquad\leq\biggl|\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})-\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|.$ (62)
For the estimation of the right-hand side of (62), we obtain
$\biggl|\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})-\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|$
$\qquad\leq\biggl|\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})-\sum\limits_{k=1}^{s}a_{k}(\lambda_{n})\biggr|+\biggl|\sum\limits_{k=1}^{s}a_{k}(\lambda_{n})-\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|$
$\qquad\leq\sum\limits_{k=s+1}^{\infty}|a_{k}(\lambda_{n})|+\sum\limits_{k=1}^{s}|a_{k}(\lambda_{n})-a_{r,k,n}(\lambda_{n})|.$
Using the estimations for $\sum\limits_{k=s+1}^{\infty}|a_{k}(\lambda_{n})|$ and $\sum\limits_{k=1}^{s}|a_{k}(\lambda_{n})-a_{r,k,n}(\lambda_{n})|$, by considering the greatest summands of them, we obtain
$\biggl|\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})-\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|<\sum\limits_{k=s+1}^{\infty}\frac{(b-a)^{k+1}}{2M\sqrt{2}^{k}\pi^{k+1}(2n-1-M)^{k}}+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M]}$
$\qquad=\frac{(b-a)^{s+2}}{2M\sqrt{2}^{s}\pi^{s+1}(2n-1-M)^{s}[\sqrt{2}\pi(2n-1-M)-(b-a)]}+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M]}$ (63)
Thus, by (62) and (63), we obtain
$|K_{n}(\lambda_{n})-K_{n}(\rho_{n})|<\frac{(b-a)^{s+2}}{2M\sqrt{2}^{s}\pi^{s+1}(2n-1-M)^{s}[\sqrt{2}\pi(2n-1-M)-(b-a)]}$
$\qquad+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M]}.$ (64)
To apply the mean value theorem, we estimate $|K_{n}^{\prime}(\lambda)|$:
$|K_{n}^{\prime}(\lambda)|=|1-g_{n}^{\prime}(\lambda)|\geq|1-|g_{n}^{\prime}(\lambda)||\geq 1-L_{n}.$ (65)
By the mean value formula, (60), (61), (64) and (65), we obtain
$|K_{n}(\lambda_{n})-K_{n}(\rho_{n})|=|K_{n}^{\prime}(\xi)||\lambda_{n}-\rho_{n}|,\qquad\xi\in[n^{2}-M,n^{2}+M]$
and
$|\lambda_{n}-\rho_{n}|=\frac{|K_{n}(\lambda_{n})-K_{n}(\rho_{n})|}{|K_{n}^{\prime}(\xi)|}<\frac{(b-a)^{s+2}}{2M\sqrt{2}^{s}\pi^{s+1}(2n-1-M)^{s}[\sqrt{2}\pi(2n-1-M)-(b-a)](1-L_{n})}$
$\qquad+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M](1-L_{n})},$
which completes the proof.
Now let us approximate $\rho_{n}$ by the fixed point iterations:
$x_{n,i+1}=n^{2}+g_{n}(x_{n,i}),$ (66)
where $g_{n}(x)$ is defined by (56). Since
$|g_{n}(\lambda_{n})|=\biggl|-C_{2n}+\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|$
$\qquad\leq|C_{2n}|+\sum\limits_{k=1}^{s}|a_{r,k,n}(\lambda_{n})|,$ (67)
first let us estimate $|a_{r,k,n}(\lambda)|$:
$\sum\limits_{k=1}^{s}|a_{r,k,n}(\lambda_{n})|<\frac{(b-a)^{2}}{M\pi^{2}(2n-1-M)}+\sum\limits_{k=2}^{r}\frac{(b-a)^{k+1}}{2M\sqrt{2}^{k}\pi^{k+1}(2n-1-M)^{k}}$
$\qquad<\frac{4}{\pi^{2}}+\sum\limits_{k=2}^{r}\frac{\sqrt{2}^{k}}{\pi^{k+1}}<\frac{4}{\pi^{2}}+\frac{2}{\pi^{2}(\pi-\sqrt{2})}.$ (68)
Therefore,
$|g_{n}(\lambda_{n})|\leq|C_{2n}|+\sum\limits_{k=1}^{s}|a_{r,k,n}(\lambda_{n})|<\frac{(b-a)}{2\pi n}+\frac{4}{\pi^{2}}+\frac{2}{\pi^{2}(\pi-\sqrt{2})}.$
On the other hand, writing (2n1)(2n-1) instead of (2n1)M(2n-1)-M in (67) and (68), we obtain
$|g_{n}(n^{2})|\leq|C_{2n}|+\sum\limits_{k=1}^{s}|a_{r,k,n}(n^{2})|$
$\qquad<\frac{(b-a)}{2\pi n}+\frac{(b-a)^{2}}{M\pi^{2}(2n-1)}+\sum\limits_{k=2}^{r}\frac{(b-a)^{k+1}}{2M\sqrt{2}^{k}\pi^{k+1}(2n-1)^{k}}$
$\qquad<\frac{(b-a)}{2\pi n}+\frac{(b-a)^{2}}{M\pi^{2}(2n-1)}+\frac{(b-a)^{3}}{2\sqrt{2}M\pi^{2}(2n-1)[\sqrt{2}\pi(2n-1)-(b-a)]},$ (69)
since $|n^{2}-k^{2}|\geq 2n-1$, for $n=1,2,\ldots$. Now we state the following result.
Theorem 5
If the assumptions about $M$ given in Condition 1 hold, then the following estimation holds for the sequence $\{x_{n,i}\}$ defined by (66):
$|x_{n,i}-\rho_{n}|<(L_{n})^{i}\biggl(\frac{(b-a)}{2\pi n(1-L_{n})}+\frac{(b-a)^{2}}{M\pi^{2}(2n-1)(1-L_{n})}$
$\qquad+\frac{(b-a)^{3}}{2\sqrt{2}M\pi^{2}(2n-1)[\sqrt{2}\pi(2n-1)-(b-a)](1-L_{n})}\biggr),$ (70)
for $i=1,2,3,\ldots$, where $L_{n}$ is defined in (58) in Theorem 4.
Proof. Without loss of generality we can take $x_{n,0}=n^{2}$. By (57), (58), and (66) we have
$|x_{n,i}-\rho_{n}|=|n^{2}+g_{n}(x_{n,i-1})-(n^{2}+g_{n}(\rho_{n}))|$
$\qquad=|g_{n}(x_{n,i-1})-g_{n}(\rho_{n})|<L_{n}|x_{n,i-1}-\rho_{n}|<(L_{n})^{i}|x_{n,0}-\rho_{n}|.$ (71)
Therefore it is enough to estimate $|x_{n,0}-\rho_{n}|$. By definitions of $\rho_{n}$ and $x_{n,0}$ we obtain
$\rho_{n}-x_{n,0}=g_{n}(\rho_{n})+n^{2}-x_{n,0}=g_{n}(\rho_{n})-g_{n}(x_{n,0})+g_{n}(n^{2})$
and by the mean value theorem there exists $x\in[n^{2}-M,n^{2}+M]$ such that
$g_{n}(\rho_{n})-g_{n}(x_{n,0})=g_{n}^{\prime}(x)(\rho_{n}-x_{n,0}).$
The last two equalities imply that
$(\rho_{n}-x_{n,0})(1-g_{n}^{\prime}(x))=g_{n}(n^{2}).$
Hence by (61), (69), and (71), we have
$|\rho_{n}-x_{n,0}|\leq\frac{|g_{n}(n^{2}))|}{1-L_{n}}<\frac{(b-a)}{2\pi n(1-L_{n})}+\frac{(b-a)^{2}}{M\pi^{2}(2n-1)(1-L_{n})}$
$\qquad+\frac{(b-a)^{3}}{2\sqrt{2}M\pi^{2}(2n-1)[\sqrt{2}\pi(2n-1)-(b-a)](1-L_{n})}$
and
$|x_{n,i}-\rho_{n}|<(L_{n})^{i}\biggl(\frac{(b-a)}{2\pi n(1-L_{n})}+\frac{(b-a)^{2}}{M\pi^{2}(2n-1)(1-L_{n})}$
$\qquad+\frac{(b-a)^{3}}{2\sqrt{2}M\pi^{2}(2n-1)[\sqrt{2}\pi(2n-1)-(b-a)](1-L_{n})}\biggr).$
The theorem is proved.
Thus by (59) and (70), we have the approximation $x_{n,i}$ for the Dirichlet eigenvalue $\lambda_{n}$ with the error
$|\lambda_{n}-x_{n,i}|<\frac{(b-a)^{s+2}}{2M\sqrt{2}^{s}\pi^{s+1}(2n-1-M)^{s}[\sqrt{2}\pi(2n-1-M)-(b-a)](1-L_{n})}$
$\qquad+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M](1-L_{n})}+(L_{n})^{i}\biggl(\frac{(b-a)}{2\pi n(1-L_{n})}$
$\qquad+\frac{(b-a)^{2}}{M\pi^{2}(2n-1)(1-L_{n})}+\frac{(b-a)^{3}}{2\sqrt{2}M\pi^{2}(2n-1)[\sqrt{2}\pi(2n-1)-(b-a)](1-L_{n})}\biggr).$
By this error formula, it is clear that the error gets smaller as $s$ and $r$ increase.
Now, we present a numerical example:
Example 1
For a=1/2a=-1/2, $b=1/2$, and $c=\pi/2$, we have the following approximations for the first Dirichlet eigenvalues $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$, $\lambda_{5}$ and $\lambda_{6}$. In our calculations, we take $s=r=5$.
$\lambda_{1}$ $=0.984205232093$
$\lambda_{2}$ $=4.003110832419$
$\lambda_{3}$ $=9.013023482675$
$\lambda_{4}$ $=16.018415152245$
$\lambda_{5}$ $=25.010308870985$
$\lambda_{6}$ $=36.010794654577.$
Usually it takes 8108-10 iterations with the tolerance 1e181e-18 by the fixed point iteration method, even if we choose an initial value that is not too close to the exact value, which means that convergence is quite fast.
Declarations
  • Funding
    Not applicable.
  • Conflict of interest
    The author states that there is no conflict of interest.
  • Ethics approval
    Not applicable.
  • Consent to participate
    Not applicable.
  • Consent for publication
    Not applicable.
  • Availability of data and materials
    Not applicable.

### v2

@@
 Abstract
-In this paper, first, we improve the asymptotic formulas obtained in [13] and obtain sharp asymptotic formulas explicitly expressed by the potential.
+In this paper, we first improve some asymptotic formulas previously obtained and provide sharp asymptotic formulas explicitly expressed by the potential.
 For the potentials of bounded variation, we obtain asymptotic formulas in which the first and second terms are explicitly determined and separated from the error terms.
@@
 We then provide estimates for the small Dirichlet eigenvalues of the one-dimensional Schrödinger operator in the Kronig-Penney model.
-Using Rouché’s theorem, we derive several useful equations from certain iteration formulas for computing these Dirichlet eigenvalues, and we estimate the eigenvalues numerically.
+We derive several useful equations from certain iteration formulas for computing these Dirichlet eigenvalues, and prove that all the eigenvalues can be found by the fixed point iteration.
+Then, using the Banach fixed point theorem, we estimate the eigenvalues numerically.
 Moreover, we present error estimates and include a numerical example.
@@
 1 Introduction and Preliminary Facts
-This paper is a continuation and contains some applications of the paper [13]. In [13], we considered the operator generated in $L_{2}[0,1]$ by the expression
-$y^{\prime\prime}+q(x)y,$ (1)
-and the boundary conditions
-$y(1)=y(0)=0,$
-where $q(x)$ is a complex-valued summable function.
-In the present paper, for brevity of some calculation, $L_{2}[0,1]$ is replaced by $L_{2}[0,\pi]$ and it is assumed that ${\displaystyle\int_{0}^{\pi}}q(x)dx=0$.
-Therefore, we consider the operator $L(q)$ generated in $L_{2}[0,\pi]$ by the expression (1) and the boundary conditions
-$y(\pi)=y(0)=0,$
-where $q(x)$ is a complex-valued summable function.
+In the present paper, we consider the operator generated in $L_{2}[0,\pi]$ by the expression
+$y^{\prime\prime}+q(x)y$ (1)
+and the boundary conditions $y(\pi)=y(0)=0$, where $q(x)$ is a complex-valued summable function.
 In [13], we proved that the eigenvalue $\lambda_{n}$ of the operator $L(q)$ lying in $O(1)$ neighborhood of $n^{2}$ satisfies the equality
-$(\lambda_{n}-n^{2}-A_{m}(\lambda_{n}))(\Psi_{n}(x),\sin nx)=R_{m+1},$ (2)
-where $n$ is a large number, $\Psi_{n}(x)$ is an eigenfunction corresponding to the eigenvalue $\lambda_{n}$ and satisfying the inequality $|(\Psi_{n},\sin nx)|>1/2$,
+$(\lambda_{n}-n^{2}-A_{m}(\lambda_{n}))(\Psi_{n},\sin nx)=R_{m+1},$ (2)
+where $n$ is a large number, $\Psi_{n}$ is an eigenfunction corresponding to the eigenvalue $\lambda_{n}$ and satisfying the inequality $|(\Psi_{n},\sin nx)|>1/2$,
 $A_{m}(\lambda_{n})=\sum_{k=0}^{m}a_{k}(\lambda_{n}),$
@@
 $R_{m+1}(\lambda_{n})=\sum_{n_{1},n_{2},...,n_{m+1}=-\infty}^{\infty}\frac{C_{n_{1}}C_{n_{2}}...C_{n_{m+1}}\bigl(q(x)\Psi_{n}(x),\sin\bigl(n+\sum_{k=1}^{m+1}n_{k}\bigr)\bigr)}{\prod_{j=1}^{m+1}[\lambda_{n}-(n+\sum_{k=1}^{j}n_{k})^{2}]},$ (6)
-and $C_{n}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos nxdx$. Here the sums for $a_{k}(\lambda_{n})$ and $R_{m+1}(\lambda_{n})$ are taken under the conditions
- $\text{ }\sum\limits_{j=1}^{s}n_{j}\neq 0,-2n,$
-for s=1,2,,ks=1,2,...,k and s=1,2,,m+1s=1,2,...,m+1, respectively. Moreover, we have
+$C_{n}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos nxdx$, and, without loss of generality, it is assumed that $C_{0}=0$.
+Here the sums for $a_{k}(\lambda_{n})$ and $R_{m+1}(\lambda_{n})$ are taken under the conditions $n_{1}+n_{2}+...+n_{s}\neq 0,-2n$, for s=1,2,,ks=1,2,...,k and s=1,2,,m+1s=1,2,...,m+1, respectively.
+Moreover, we have
 $a_{k}=O\biggl(\bigl(\frac{ln\left|n\right|}{n}\bigr)^{k}\biggr),\qquad R_{m+1}=O\biggl(\bigl(\frac{ln\left|n\right|}{n}\bigr)^{m+1}\biggr).$ (7)
@@
 In this paper, in Section 2, we improve the asymptotic formulas, instead of the series in (8), we write some Fourier coefficients of some functions explicitly expressed by the potential $q$.
-For the potential of bounded variation, we explicitly determine the first and second terms of the asymptotic formulas, which have the order $O(n^{-1})$ and $O(n^{-2})$, respectively (see (33) and Theorem 1).
-Note that, generally speaking, if the potential $q$ is a piecewise continuous function with some jump, then the first term $C_{2n}$ is of order $n^{-1}$ and the next term in the asymptotic formula (33) is of order $n^{-2}$.
-Consequently, asymptotic formula (33) with error $O(n^{-3})$ explicitly separates the first and second terms from errors.
+For the potential of bounded variation, we explicitly determine the first and second terms of the asymptotic formulas, which have the orders $O(n^{-1})$ and $O(n^{-2})$, respectively (see (30) and Theorem 1).
+Note that, generally speaking, if the potential $q$ is a piecewise continuous function with some jump, then the first term $C_{2n}$ is of order $n^{-1}$ and the next term in asymptotic formula (30) is of order $n^{-2}$.
+Consequently, asymptotic formula (30) with error $O(n^{-3})$ explicitly separates the first and second terms from the error terms.
 Then, we demonstrate this formula for the Kronig-Penney potential
 $q(x)=\begin{cases}a&\text{if }x\in[0,c]\\[5.69054pt] b&\text{if }x\in(c,\pi],\end{cases}$ (9)
-where $q(x+\pi)=q(x)$, and $c\in(0,\pi)$. Without loss of generality, we assume that $a<b$, and
-$q_{0}=\frac{1}{\pi}\int_{0}^{\pi}q(x)dx=0.$
-Therefore, we have
-$ac+b(\pi-c)=0,\qquad a<0<b,$ (10)
-and $(b-a)c=b\pi$.
+where $q(x+\pi)=q(x)$, and $c\in(0,\pi)$. Without loss of generality, we assume that $a<b$. This inequality with the condition $C_{0}=0$ implies that
+$ac+b(\pi-c)=0,\qquad a<0<b,\text{ }(b-a)c=b\pi.$ (10)
 Note that, the Kronig-Penney model is a simplified model for an electron in a one-dimensional periodic potential and has been studied in many works (see, for example, [1, 3, 8, 10, 11], and references therein).
-Veliev [10, 11] studied the bands and gaps in the spectrum of the Schrödinger operator with Kronig-Penney potential and obtained asymptotic formulas for the length of the gaps in the spectrum.
+Veliev [10, 11] studied the bands and gaps in the spectrum of the Schrödinger operator with the Kronig-Penney potential and obtained asymptotic formulas for the length of the gaps in the spectrum.
 In Section 3, we provide estimates for small Dirichlet eigenvalues of the Schrödinger operator $L(q)$ with the Kronig-Penney potential.
-We obtain some useful equations for calculating the Dirichlet eigenvalues using Rouche’s theorem.
-These equations are derived from some iterations formulas by the methods used in [7, 9, 12, 13].
-It is important to note that in these papers, the authors used asymptotic formulas for large eigenvalues, which cannot be applied to small eigenvalues.
-Therefore, in Section 3, we examine the small eigenvalues using numerical methods.
-Moreover, as a result of meticulous and detailed investigations, we establish conditions on the potential under which all small eigenvalues satisfy the equations obtained from the iterative formulas.
+First, we derive equation (42) for calculating the Dirichlet eigenvalues using some iterations formulas by the methods used in [7, 9, 12, 13].
+It is important to note that, the main difficulty of this section is to prove that all the eigenvalues (small and large) of $L(q)$ can be obtained from (42) by the fixed point iteration.
+Then, we examine the small eigenvalues using numerical methods.
+Moreover, in Section 3, as a result of meticulous and detailed investigations, we establish conditions on the potential under which all the small eigenvalues satisfy equality (42) and the Banach fixed point theorem can be used for this equality.
 2 On the Sharp Asymptotic Formulas
 In this section, first, to obtain explicit formulas for the large eigenvalues from (8), we consider the series in (8) as the difference of the series
-$A_{1,n}(n^{2}):=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}^{2}}{n^{2}-(n+k)^{2}}$ (11)
-and
-$B_{1,n}(n^{2})=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}C_{k+2n}}{n^{2}-(n+k)^{2}}.$ (12)
-Then, we consider these series in the following lemmas. First, let us consider $A_{1,n}(n^{2})$. For this introduce the notations:
+$A_{1,n}(n^{2}):=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}^{2}}{n^{2}-(n+k)^{2}},\qquad B_{1,n}(n^{2}):=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}C_{k+2n}}{n^{2}-(n+k)^{2}}.$ (11)
+Then, we consider these series in the following lemmas. First, let us consider $A_{1,n}(n^{2})$. For this, introduce the notations:
 $Q(x,n)=\int_{0}^{x}q(t)e^{i2nt}\,dt-q_{-2n}x,\qquad Q_{n,0}=\frac{1}{\pi}\int_{0}^{\pi}Q(x,n)dx$ (12)
@@
 $G(x,n)=\int_{0}^{x}\bigl(q(t)e^{i2nt}\,+\frac{i\pi}{2}e^{it}\,q_{-2n}\bigr)dt,$ (13)
-where
-$q_{k}=\frac{1}{\pi}\int_{0}^{\pi}q(x)e^{-ikx}dx\,,\qquad q_{0}=0.$
+where $q_{k}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)e^{-ikx}dx$ and $q_{0}=0$.
 Lemma 1
@@
 where $D(n,f)=D_{1}(n,f)+D_{2}(n,f)$,
-$D_{1}(n,f)=:\frac{i}{4n\pi}\int_{0}^{\pi}f(x)(Q(x,n)-Q_{n,0})e^{-i2nx}\,dx,\text{ }$
-$D_{2}(n,f)=:\frac{i}{4n\pi}\int_{0}^{\pi}f(x)G(x,n)e^{-i2nx}\,dx,$
+$D_{1}(n,f)=\frac{i}{4n\pi}\int_{0}^{\pi}f(x)(Q(x,n)-Q_{n,0})e^{-i2nx}\,dx,\quad D_{2}(n,f)=\frac{i}{4n\pi}\int_{0}^{\pi}f(x)G(x,n)e^{-i2nx}\,dx,$
 and p(x)=q(x)p(x)=q(-x).
@@
 $A_{1,n}(n^{2})=-\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}^{2}}{k(2n+k)}.$
-Since
-$\frac{1}{k(2n+k)}=\frac{1}{2n}\biggl(\frac{1}{k}-\frac{1}{2n+k}\biggr),$
-we see that
+Since $\dfrac{1}{k(2n+k)}=\dfrac{1}{2n}\bigl(\dfrac{1}{k}-\dfrac{1}{2n+k}\bigr)$, we see that
 $A_{1,n}(n^{2})=\frac{1}{2n}\sum_{k\neq 0,-2n}\frac{C_{k}^{2}}{2n+k}-\frac{1}{2n}\sum_{k\neq 0,-2n}\frac{C_{k}^{2}}{k}.$ (15)
@@
 $\frac{1}{2n}\sum_{k\neq 0,-2n}\frac{C_{k}^{2}}{k}=\frac{C_{2n}^{2}}{4n^{2}}.$ (16)
-Now let us estimate the first term on the right side of (16). Since $q_{-k}=\overline{q_{k}}$ and $\cos kx=\frac{1}{2}(e^{ikx}+e^{-ikx})$ we have
-$C_{k}^{2}=\frac{1}{4}(2\left|q_{k}\right|^{2}+q_{k}^{2}+q_{-k}^{2}).$
+Now let us estimate the first term on the right side of (15).
+Since $q_{-k}=\overline{q_{k}}$ and $\cos kx=\dfrac{1}{2}(e^{ikx}+e^{-ikx})$, we have $C_{k}^{2}=\dfrac{1}{4}(2\left|q_{k}\right|^{2}+q_{k}^{2}+q_{-k}^{2})$.
 Therefore, the first term on the right side of (15) can be written in the form
@@
 $I_{1,1}=D_{1}(n,q),\qquad I_{1,2}=D_{2}(n,q),\qquad I_{1}=D(n,q).$ (19)
-It is clear that
-$Q(\pi,n)=\int_{0}^{\pi}q(t)e^{i2nt}\,dt-q_{-2n}\pi=0,\qquad Q(0,n)=0$
-and that the derivative of $Q(x,n)$ with respect to $x$ is
-$Q^{\prime}(x,n)=q(x)e^{i2nx}\,-q_{-2n}.$
-Therefore, using the integration by parts, we see that
+It is clear that $Q(\pi,n)=0$, $Q(0,n)=0$, and that the derivative of $Q(x,n)$ with respect to $x$ is $Q^{\prime}(x,n)=q(x)e^{i2nx}\,-q_{-2n}$. Using the integration by parts, we see that
 $\int_{0}^{\pi}Q(x,n)e^{-i(2n+2m)x}dx=\frac{\pi q_{2m}}{i(2n+2m)},$
@@
 $G(\pi,n)=\int_{0}^{\pi}\bigl(q(t)e^{i2nt}\,+\frac{i\pi}{2}e^{it}\,q_{-2n}\bigr)dt=\pi q_{-2n}+\frac{\pi}{2}(e^{i\pi}-1)q_{-2n}=0,$
-$G(0,n)=0$ and
-$G^{\prime}(x,n)=q(x)e^{i2nx}\,+\frac{i\pi}{2}e^{ix}q_{-2n}.$
-Therefore,
+$G(0,n)=0$ and that $G^{\prime}(x,n)=q(x)e^{i2nx}\,+\dfrac{i\pi}{2}e^{ix}q_{-2n}$. Therefore,
 $\int_{0}^{\pi}G(x,n)e^{-i(2n+2m+1)x}dx=\frac{\pi q_{2m+1}}{i(2n+2m+1)}.$
@@
 $I_{2}=D(n,p),\qquad I_{2}+I_{3}=2\operatorname{Re}D(n,p).$ (20)
-Thus the proof of (15) follows from (16)-(21). The lemma is proved.
+Thus, the proof of (14) follows from (15)-(20). The lemma is proved.
 Now let us consider $B_{1,n}(n^{2})$.
@@
 Proof.
-To prove (22), we estimate the left and write side of this equality separately.
-First estimate the right-side.
+To prove (21), we estimate the left and right sides of this equality separately.
+First, estimate the right-side.
 Let $Q_{k}={\displaystyle\int_{0}^{\pi}}Q(x)\sin kxdx$.
@@
 of $Q(x)$.
-Using the equality $(Q^{2}(x))^{\prime}=2q(x)Q(x)$, the decomposition of $Q(x)$, and integration by parts, and then the equality $\sin kx\sin 2nx=\frac{1}{2}(\cos(2n-k)x-\cos(2n+k)x)$, we obtain
+Using the equality $(Q^{2}(x))^{\prime}=2q(x)Q(x)$, integration by parts, the decomposition of $Q(x)$, and then the equality $\sin kx\sin 2nx=\dfrac{1}{2}(\cos(2n-k)x-\cos(2n+k)x)$, we obtain
 $\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx=-\frac{1}{n}\int_{0}^{\pi}q(x)Q(x)\sin 2nxdx$
-$\qquad=-\frac{2}{\pi n}\int_{0}^{\pi}q(x)\sum_{k=1}^{\infty}Q_{k}\sin kx\sin 2nxdx$
 $\qquad=-\frac{1}{\pi n}\int_{0}^{\pi}q(x)\sum_{k=1}^{\infty}Q_{k}(\cos(2n-k)x-\cos(2n+k)x)dx$
@@
 Now, using the equalities $Q_{k}=\pi C_{k}/k$, $C_{-k}=C_{k}$, and doing the substitution $k\rightarrow-k$ in the second summation of the last equality, we obtain
-$\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx=\frac{-\pi}{n}\sum_{k\in\mathbb{Z},k\neq 0}\frac{C_{k}C_{2n-k}}{k}$
+$\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx=\frac{-\pi}{n}\sum_{k\in\mathbb{Z},k\neq 0}\frac{C_{k}C_{2n-k}}{k}.$
 Thus, the right side of (21) has the form
@@
 since $C_{0}=0$.
-Now, we prove that the left side of (22) also can be written in this form. Doing the substitution $k\rightarrow-k$, and then using the equality
-$\frac{1}{k(2n-k)}=\frac{1}{2n}\biggl(\frac{1}{k}+\frac{1}{2n-k}\biggr),$
-we see that
+Now, we prove that the left side of (21) also can be written in this form.
+Doing the substitution $k\rightarrow-k$, and then using the equality $\dfrac{1}{k(2n-k)}=\dfrac{1}{2n}\bigl(\dfrac{1}{k}+\dfrac{1}{2n-k}\bigr)$, we see that
 $B_{1,n}(n^{2})=\sum_{k\in\mathbb{Z},k\neq 0,-2n}\frac{C_{k}C_{2n+k}}{n^{2}-(n+k)^{2}}=\sum_{k\in\mathbb{Z},k\neq 0,-2n}\frac{C_{k}C_{2n+k}}{(2n+k)(-k)}=\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k(2n-k)}$
@@
 On the other hand, doing the substitution $k\rightarrow 2n-k$ in the second summation of the last equality, we obtain
-$\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{2n-k}=\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k}.$
-Therefore, we have
-$B_{1,n}(n^{2})=\frac{1}{n}\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k},$
+$\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{2n-k}=\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k},\qquad B_{1,n}(n^{2})=\frac{1}{n}\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k},$
 which means that (21) holds. The lemma is proved.
 Now, we consider $a_{2}(n^{2})$. It is clear that $a_{2}(n^{2})=a_{2,1}(n^{2})-a_{2,2}(n^{2})$, where
-$a_{2,1}(n^{2})=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{k+l}}{[n^{2}-(n+k)^{2}][n^{2}-((n+k+l))^{2}]}$
-$\qquad=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(2n+k)(k+l)(2n+k+l)}$
-and
+$a_{2,1}(n^{2})$ $=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{k+l}}{[n^{2}-(n+k)^{2}][n^{2}-((n+k+l))^{2}]}=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(2n+k)(k+l)(2n+k+l)},$
 $a_{2,2}(n^{2})$ $=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(2n+k)(k+l)(2n+k+l)}.$
-Using the equality
-$\frac{1}{(k+l)(2n+k+l)}=\frac{1}{2n}\left(\frac{1}{k+l}+\frac{1}{2n+k+l}\right),$
-for $l=0$ and $l\neq 0$, we obtain
+Using the equality $\dfrac{1}{(k+l)(2n+k+l)}=\dfrac{1}{2n}\bigl(\dfrac{1}{k+l}+\dfrac{1}{2n+k+l}\bigr)$, for $l=0$ and $l\neq 0$, we obtain
 $a_{2,1}(n^{2})=\frac{1}{4n^{2}}(S_{1}(n,q)+S_{2}(n,q)+S_{3}(n,q)+S_{4}(n,q)),$ (22)
 where
-$S_{1}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(k+l)},$
-$S_{2}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{(2n+k)(k+l)},$
-$S_{3}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(2n+k+l)},$
-and
-$S_{4}(n,q)=\sum\limits_{k,k+l\neq 0,2n}\frac{C_{k}C_{l}C_{-k-l}}{(2n+k)(2n+k+l)}.$
+$S_{1}(n,q)$ $=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(k+l)},\qquad S_{2}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{(2n+k)(k+l)},$
+$S_{3}(n,q)$ $=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(2n+k+l)},\qquad S_{4}(n,q)=\sum\limits_{k,k+l\neq 0,2n}\frac{C_{k}C_{l}C_{-k-l}}{(2n+k)(2n+k+l)}.$
 First, let us consider $S_{1}(n,q)$. Grouping the terms
@@
 $\frac{C_{l}C_{-k-l}C_{k}}{l(-k)},\quad\frac{C_{-k-l}C_{l}C_{k}}{(-k-l)(-k)},\quad\frac{C_{-k-l}C_{k}C_{l}}{(-k-l)(-l)}$
-with the equal multiplicands,
-$C_{k}C_{l}C_{-k-l},\quad C_{l}C_{k}C_{-k-l},\quad C_{k}C_{-k-l}C_{l},$
-$C_{l}C_{-k-l}C_{k},\quad C_{-k-l}C_{l}C_{k},\quad C_{-k-l}C_{k}q_{l}$
-and using the equality
-$\frac{1}{k(k+l)}+\frac{1}{l(k+l)}=\frac{1}{kl},$
-we obtain that
+with the equal multiplicands, $C_{k}C_{l}C_{-k-l}$, $C_{l}C_{k}C_{-k-l}$, $C_{k}C_{-k-l}C_{l}$, $C_{l}C_{-k-l}C_{k}$, $C_{-k-l}C_{l}C_{k}$, $C_{-k-l}C_{k}C_{l}$ and using the equality $\dfrac{1}{k(k+l)}+\dfrac{1}{l(k+l)}=\dfrac{1}{kl}$, we obtain that
 $\sum\limits_{k,l}\frac{C_{k}C_{l}C_{-k-l}}{k(k+l)}=0,$ (23)
@@
 $S_{1,1}(n,q):=\frac{C_{2n}}{2n}O(1)=o\bigl(\frac{1}{n}\bigr),\qquad S_{1,2}(q):=\frac{C_{2n}}{2n}O(1)=o\bigl(\frac{1}{n}\bigr),$ (24)
-for $q\in L_{2}[0,1]$.
+for $q\in L_{2}[0,\pi]$.
 Denote by $E$ the set of all functions satisfying
-$C_{n}=\int_{0}^{\pi}q(x)\cos nxdx=O\bigl(\frac{1}{n}\bigr)$ (26)
+$C_{n}=\frac{1}{\pi}\int_{0}^{\pi}q(x)\cos nxdx=O\bigl(\frac{1}{n}\bigr),$ (25)
 as $n\rightarrow\infty$.
@@
 Here we do not discuss the set $E$, since this is not the aim of this paper.
-We will apply the obtained further result for potential (9) which is a bounded variation.
+We will apply the obtained further results for potential (9) which is a bounded variation.
 Using (25) in (24), we obtain
@@
 If $q\in E$, then $S_{1}(n,q)=O(n^{-2})$.
-Now, we estimate $S_{j}(n,q)$, for $j=2,3,4$, in the following lemma. For this, we use the following easily checkable relations
-$\sum\limits_{k\neq 0,-2n}\frac{1}{k^{2}(2n+k)}=O\bigl(\frac{1}{n}\bigr),$ (27)
-$\sum\limits_{l\neq 0,-k}\frac{1}{l(k+l)^{2}}=O\bigl(\frac{1}{k}\bigr),$ (28)
-and
-$\sum\limits_{k\neq 0,-2n}\frac{1}{\left|k(2n+k)\right|}=O\bigl(\frac{\ln n}{n}\bigr).$ (29)
-To prove (27), split the sum into three parts
+Now, we estimate $S_{j}(n,q)$, for $j=2,3,4$, in the following lemma. For this, we use the following easily checkable relations:
+$\sum\limits_{k\neq 0,-2n}\frac{1}{k^{2}(2n+k)}=O\bigl(\frac{1}{n}\bigr),\quad\sum\limits_{l\neq 0,-k}\frac{1}{l(k+l)^{2}}=O\bigl(\frac{1}{k}\bigr),\quad\sum\limits_{k\neq 0,-2n}\frac{1}{\left|k(2n+k)\right|}=O\bigl(\frac{\ln n}{n}\bigr).$ (26)
+To prove the first equality, split the sum into three parts
 $\sum\limits_{k\leq n,\text{ }k\neq 0}\frac{1}{k^{2}(2n+k)},\qquad\sum\limits_{k>n}\frac{1}{k^{2}(2n+k)},\qquad\sum\limits_{k<-n,k\neq-2n}\frac{1}{k^{2}(2n+k)}$
-and estimate each of them in a standard way. Equalities (28) and (29) can be proved in the same way.
+and estimate each of them in a standard way. The second and third equalities can be proved in the same way.
 Lemma 3
@@
 $\left|-S_{2}(n,q)\right|=O(1)\biggl(\sum\limits_{k\neq 0,-2n}\sum\limits_{l\neq 0,-k}\frac{1}{\left|k(2n+k)l(k+l)^{2}\right|}\biggr).$
-Now using (28) and then (27), we obtain
+Now using (26), we obtain
 $\left|-S_{2}(n,q)\right|=O(1)\sum\limits_{k\neq 0,-2n}\frac{1}{\left|k^{2}(2n+k)\right|}=O\bigl(\frac{1}{n}\bigr).$
@@
 $\left|-S_{3}(n,q)\right|=O(1)\sum\limits_{s\neq 0,-2n}\sum\limits_{k\neq 0,s}\frac{1}{\left|s(2n+s)k^{2}(s-k)\right|}.$
-Now using (27) and (28), we obtain
+Now using (26), we obtain
 $\left|-S_{3}(n,q)\right|=O(1)\sum\limits_{s\neq 0,2n}\frac{1}{\left|s^{2}(2n+s)\right|}=O\bigl(\frac{1}{n}\bigr),$
@@
 $a_{2,1}(n^{2})=O\bigl(\frac{1}{n^{3}}\bigr),$ (28)
-if $q\in E$. Now let us estimate $a_{2,2}(n^{2})$. Arguing as in the proof of (23), we see that
-$a_{2,2}(n^{2})=\frac{1}{4n^{2}}(H_{1}+H_{2}+H_{3}+H_{4}),$
-where
-$H_{1}(q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(k+l)},$
-$H_{2}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{(2n+k)(k+l)},$
-$H_{3}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(2n+k+l)},$
-and
-$H_{4}(n,q)=\sum\limits_{k,k+l\neq 0,2n}\frac{C_{k}C_{l}C_{2n+k+l}}{(2n+k)(2n+k+l)}.$
+if $q\in E$. Now, let us estimate $a_{2,2}(n^{2})$. Arguing as in the proof of (22), we see that $a_{2,2}(n^{2})=\dfrac{1}{4n^{2}}(H_{1}+H_{2}+H_{3}+H_{4})$, where
+$H_{1}(n,q)$ $=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(k+l)},\qquad H_{2}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{(2n+k)(k+l)},$
+$H_{3}(n,q)$ $=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(2n+k+l)},\qquad H_{4}(n,q)=\sum\limits_{k,k+l\neq 0,2n}\frac{C_{k}C_{l}C_{2n+k+l}}{(2n+k)(2n+k+l)}.$
 Therefore, arguing as in the proof of Lemma 3, we conclude that
@@
 If $q\in E$, then the following asymptotic formula holds:
-$\lambda_{n}=n^{2}-C_{2n}-\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)+O\bigl(\frac{1}{n^{3}}\bigr),$ (33)
+$\lambda_{n}=n^{2}-C_{2n}+\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)+O\bigl(\frac{1}{n^{3}}\bigr),$ (30)
 where $D(n,f)$ and $Q(x)$ are defined in Lemma 1 and Lemma 2.
-Proof. It follows from (8), (26) and (27) that
-$\lambda_{n}=n^{2}+O\bigl(\frac{1}{n}\bigr).$
-Using this, we easily get the estimate
-$\sum_{\begin{subarray}{c}k\neq 0,-2n\end{subarray}}^{\infty}\mid\frac{1}{\lambda_{n}-(n+k)^{2}}-\frac{1}{n^{2}-(n+k)^{2}}\mid=O\bigl(\frac{1}{n^{3}}\bigr).$
+Proof. It follows from (8), (25) and (26) that $\lambda_{n}=n^{2}+O\bigl(\dfrac{1}{n}\bigr)$. Using this, we easily get the estimate
+$\sum_{\begin{subarray}{c}k\neq 0,-2n\end{subarray}}^{\infty}\biggl|\frac{1}{\lambda_{n}-(n+k)^{2}}-\frac{1}{n^{2}-(n+k)^{2}}\biggr|=O\bigl(\frac{1}{n^{3}}\bigr).$
 By virtue of this estimate
 $a_{j}\left(\lambda_{n}\right)=a_{j}\left(n^{2}\right)+O\bigl(\frac{1}{n^{3}}\bigr),$
-for $j=1,2,\ldots$. Therefore it follows from (2)-(6) that
+for $j=1,2,\ldots$. Therefore, it follows from (2)-(6) that
 $\lambda_{n}=n^{2}-C_{2n}+a_{1}\left(n^{2}\right)+a_{2}\left(n^{2}\right)+a_{3}\left(n^{2}\right)+O\bigl(\frac{1}{n^{3}}\bigr),$ (31)
-where
-$a_{1}\left(n^{2}\right)=D(n,q)+2\operatorname{Re}D(n,p)+O\bigl(\frac{1}{n^{3}}\bigr)$ (35)
-(see Lemma 1, Lemma 2 and (26)),
+where $a_{1}(n^{2})=A_{1,n}(n^{2})-B_{1,n}(n^{2})$, equivalently
+$a_{1}\left(n^{2}\right)=\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)+O\bigl(\frac{1}{n^{3}}\bigr)$ (32)
+(see Lemma 1, Lemma 2 and (25)), and
 $a_{2}\left(n^{2}\right)=O\bigl(\frac{1}{n^{3}}\bigr)$ (33)
@@
 $a_{3}\left(n^{2}\right)=\sum_{n_{1},n_{2},...,n_{k}=-\infty}^{\infty}\frac{C_{k}C_{l}C_{s}(C_{k+l+s}-C_{2n+k+l+s})}{k(2n+k)(k+l)(2n+k+l)(k+l+s)(2n+k+l+s)}.$
-Moreover, arguing as in the proof of (36) we see that $a_{3}\left(n^{2}\right)=O\bigl(\dfrac{1}{n^{3}}\bigr)$. Thus, the proof of the theorem follows from (34)-(36).
+Moreover, arguing as in the proof of (33), we see that $a_{3}\left(n^{2}\right)=O\bigl(\dfrac{1}{n^{3}}\bigr)$. Thus, the proof of the theorem follows from (31)-(33).
 Now, let us demonstrate the obtained results for potential (9). In this case, by direct calculation, we have
-$q_{k}=\dfrac{1}{\pi}\int_{0}^{\pi}q(x)e^{-ikx}dx\,=\dfrac{1}{i\pi k}((b-a)e^{-ikc}+a-(-1)^{k}b).$
+$q_{k}=\dfrac{1}{\pi}\int_{0}^{\pi}q(x)e^{-ikx}dx\,=\dfrac{1}{i\pi k}((b-a)e^{-ikc}+a-(-1)^{k}b).$ (34)
 Using the equalities $Q(\pi,n)=G(\pi,n)=0$ (see the proof of Lemma 1), we obtain
@@
 $G(x,n)=\begin{cases}\dfrac{a}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)&\text{if }x\in[0,c]\\[5.69054pt] \dfrac{b}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)+\pi q_{-2n}&\text{if }x\in(c,\pi],\end{cases}$
-where $Q(x,n)$, and $G(x,n)$ are defined by (13) and (14). Therefore, by Lemma 1 we have
+where $Q(x,n)$, and $G(x,n)$ are defined by (12) and (13). Therefore, by Lemma 1, we have
 $D_{1}(n,q)=\frac{i}{4\pi n}\int_{0}^{\pi}q(x)Q(x,n)e^{-i2nx}\,dx-\frac{iQ_{n,0}}{4\pi n}\int_{0}^{\pi}q(x)e^{-i2nx}\,dx$
@@
 $\qquad=\frac{i}{4\pi n}\biggl(\int_{0}^{c}q(x)G(x,n)e^{-i2nx}\,dx+\int_{c}^{\pi}q(x)G(x,n)e^{-i2nx}\,dx\biggr).$
-We first calculate the terms in $D_{1}(n,q)$. Using $(b-a)c=b\pi$ or $ac=b(c-\pi)$ by (10) and $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$, we have
+We first calculate the terms in $D_{1}(n,q)$. Using $(b-a)c=b\pi$ or $ac=b(c-\pi)$ by (10) and substituting $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$ by (34), we have
 $\pi Q_{n,0}$ $=\int_{0}^{c}\bigl(\dfrac{a}{i2n}(e^{i2nx}-1)-q_{-2n}x\bigr)dx+\int_{c}^{\pi}\bigl(\dfrac{b}{i2n}(e^{i2nx}-1)-q_{-2n}x+\pi q_{-2n}\bigr)dx$
-$=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}+\pi(\frac{\pi}{2}-c)q_{-2n}=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}+\pi(\frac{\pi}{2}-c)\frac{i(b-a)(e^{i2nc}-1)}{2\pi n}$
-$=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}-\frac{i\pi(b+a)(e^{i2nc}-1)}{4n},$ (37)
+$=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}+\pi(\frac{\pi}{2}-c)q_{-2n}=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}-\frac{i\pi(b+a)(e^{i2nc}-1)}{4n},$
 $\frac{i}{4n}Q_{n,0}q_{2n}=\frac{i(b^{2}-a^{2})(1-\cos 2nc)}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{16\pi^{2}n^{4}},$
 $\int_{0}^{c}q(x)Q(x,n)e^{-i2nx}\,dx=\int_{0}^{c}a\biggl(\dfrac{a}{i2n}(e^{i2nx}-1)-q_{-2n}x\biggr)e^{-i2nx}\,dx$
-$\qquad=\dfrac{a^{2}}{i2n}\int_{0}^{c}(1-e^{-i2nx})\,dx-aq_{-2n}\int_{0}^{c}xe^{-i2nx}\,dx$
 $\qquad=\dfrac{a^{2}}{i2n}\biggl(c+\dfrac{e^{-i2nc}-1}{i2n}\biggr)+\dfrac{aq_{-2n}}{i2n}\biggl(ce^{-i2nc}+\dfrac{e^{-i2nc}-1}{i2n}\biggr),$
@@
 $\int_{c}^{\pi}q(x)Q(x,n)e^{-i2nx}\,dx=\int_{c}^{\pi}b\biggl(\dfrac{b}{i2n}(e^{i2nx}-1)-q_{-2n}x+\pi q_{-2n}\biggr)e^{-i2nx}\,dx$
-$\qquad=\dfrac{b^{2}}{i2n}\int_{c}^{\pi}(1-e^{-i2nx})\,dx-bq_{-2n}\int_{c}^{\pi}xe^{-i2nx}\,dx+b\pi q_{-2n}\int_{c}^{\pi}e^{-i2nx}\,dx$
 $\qquad=\dfrac{b^{2}}{i2n}\biggl(\pi-c-\dfrac{e^{-i2nc}-1}{i2n}\biggr)+\dfrac{bq_{-2n}}{i2n}\biggl(\pi-ce^{-i2nc}-\dfrac{e^{-i2nc}-1}{i2n}\biggr)+\dfrac{b\pi q_{-2n}}{i2n}(e^{-i2nc}-1).$
@@
 $\qquad=\frac{i}{4\pi n}\biggl(\frac{(b^{2}-a^{2})(e^{-i2nc}-1)}{4n^{2}}-\frac{ab\pi}{i2n}+\frac{(b-a)(e^{-i2nc}-1)q_{-2n}}{4n^{2}}\biggr)-\frac{iQ_{n,0}q_{2n}}{4n}$
-and substituting $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$ in the last expression, we have
-$D_{1}(n,q)=-\frac{ab}{8n^{2}}+\frac{i(b^{2}-a^{2})(e^{-i2nc}-1)}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{16\pi^{2}n^{4}}$
-$\qquad+\frac{i(b^{2}-a^{2})(1-\cos 2nc)}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{16\pi^{2}n^{4}}$
-$\qquad=-\frac{ab}{8n^{2}}+\frac{(b^{2}-a^{2})\sin 2nc}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{8\pi^{2}n^{4}}.$ (38)
+and by (34), we have
+$D_{1}(n,q)=-\frac{ab}{8n^{2}}+\frac{(b^{2}-a^{2})\sin 2nc}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{8\pi^{2}n^{4}}.$ (35)
 Now let us calculate the terms in $D_{2}(n,q)$:
@@
 $\qquad-\frac{\pi(b-a)(1-e^{-i2nc})q_{-2n}}{8n^{2}}-\frac{\pi q_{-2n}(b+a+(b-a)e^{-i(2n-1)c})}{2(2n-1)^{2}}\biggr).$
-Using $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$, we obtain
+Using (34), we have
 $D_{2}(n,q)=-\frac{ab}{8n^{2}}+\frac{i(b^{2}-a^{2})(e^{-i2nc}-1)}{16\pi n^{3}}-\frac{ib(b-a)(1-e^{-i2nc})(e^{i2nc}+e^{ic})}{16n^{2}(2n-1)}+\frac{ib(b-a)(\cos 2nc-1)}{16n^{3}}$
 $\qquad+\frac{(b-a)^{2}(\cos 2nc-1)}{32\pi n^{4}}+\frac{(b-a)(e^{-i2nc}-1)[(b+a)e^{i2nc}+(b-a)e^{ic}]}{16\pi n^{2}(2n-1)^{2}}.$ (36)
-Thus, by (38), (39), and Lemma 1, we write
-$D(n,q)=D_{1}(n,q)+D_{2}(n,q)=-\frac{ab}{4n^{2}}+\frac{(b^{2}-a^{2})\sin 2nc}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{8\pi^{2}n^{4}}$
-$\qquad+\frac{i(b^{2}-a^{2})(e^{-i2nc}-1)}{16\pi n^{3}}+\frac{ib(b-a)(\cos 2nc-1)}{16n^{3}}-\frac{ib(b-a)(1-e^{-i2nc})(e^{i2nc}+e^{ic})}{16n^{2}(2n-1)}$
-$\qquad+\frac{(b-a)^{2}(\cos 2nc-1)}{32\pi n^{4}}+\frac{(b-a)(e^{-i2nc}-1)[(b+a)e^{i2nc}+(b-a)e^{ic}]}{16\pi n^{2}(2n-1)^{2}}$
-or
+Therefore, using (35), (36), Lemma 1, and simplifying a bit, we obtain
 $D(n,q)=D_{1}(n,q)+D_{2}(n,q)=-\frac{ab}{4n^{2}}+\frac{(b^{2}-a^{2})\sin 2nc}{8\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{8\pi^{2}n^{4}}$
@@
 $\qquad+\frac{(b-a)^{2}(\cos 2nc-1)}{32\pi n^{4}}+\frac{(b-a)(e^{-i2nc}-1)[(b+a)e^{i2nc}+(b-a)e^{ic}]}{16\pi n^{2}(2n-1)^{2}}.$ (37)
-Now, let us calculate $\operatorname{Re}D(n,p)$, where $p(x)$ is defined in Lemma 1. For the Kronig-Penney potential, $p(x)$ becomes
-$p(x)=\begin{cases}b&\text{if }x\in[0,\pi-c),\\[5.69054pt] a&\text{if }x\in[\pi-c,\pi]\end{cases}.$
-Therefore,
+Now, let us calculate $\operatorname{Re}D(n,p)$, where $p(x)$ is defined in Lemma 1.
+For the Kronig-Penney potential, $p(x)$ becomes $p(x)=b$ for $x\in[0,\pi-c)$ and $p(x)=a$ for $x\in[\pi-c,\pi]$.
+Therefore, we have
 $D_{1}(n,p)=\frac{i}{4\pi n}\int_{0}^{\pi}p(x)Q(x,n)e^{-i2nx}\,dx-\frac{iQ_{n,0}}{4\pi n}\int_{0}^{\pi}p(x)e^{-i2nx}\,dx$
@@
 and
-$D_{2}(n,p)=\frac{i}{4\pi n}\int_{0}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx$
-$\qquad=\frac{i}{4\pi n}\biggl(\int_{0}^{\pi-c}p(x)G(x,n)e^{-i2nx}\,dx+\int_{\pi-c}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx\biggr).$
-We first calculate the terms in $D_{1}(n,p)$. Using $(b-a)c=b\pi$ or $ac=b(c-\pi)$ by (10), $p_{2n}=q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$, and (37), we have
+$D_{2}(n,p)=\frac{i}{4\pi n}\biggl(\int_{0}^{\pi-c}p(x)G(x,n)e^{-i2nx}\,dx+\int_{\pi-c}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx\biggr).$
+By direct calculations, similar to the calculations above for the terms in $D_{1}(n,q)$ and $D_{2}(n,q)$, we obtain
 $\frac{i}{4n}Q_{n,0}p_{2n}=-\frac{i(b^{2}-a^{2})(e^{i2nc}-1)^{2}}{32\pi n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{32\pi^{2}n^{4}},$
-$\int_{0}^{\pi-c}p(x)Q(x,n)e^{-i2nx}\,dx=\int_{0}^{\pi-c}b\biggl(\dfrac{a}{i2n}(e^{i2nx}-1)-q_{-2n}x\biggr)e^{-i2nx}\,dx$
-$\qquad=\dfrac{ab}{i2n}\int_{0}^{\pi-c}(1-e^{-i2nx})\,dx-bq_{-2n}\int_{0}^{\pi-c}xe^{-i2nx}\,dx$
-$\qquad=\dfrac{ab}{i2n}\biggl(\pi-c+\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{bq_{-2n}}{i2n}\biggl((\pi-c)e^{i2nc}+\dfrac{e^{i2nc}-1}{i2n}\biggr),$
+$D_{1}(n,p)=\frac{ab}{8n^{2}}-\frac{i(b-a)(b+3a)(e^{i2nc}-1)^{2}}{32\pi n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{16\pi^{2}n^{4}},$ (38)
 and
-$\int_{\pi-c}^{\pi}p(x)Q(x,n)e^{-i2nx}\,dx=\int_{\pi-c}^{\pi}a\biggl(\dfrac{b}{i2n}(e^{i2nx}-1)-q_{-2n}x+\pi q_{-2n}\biggr)e^{-i2nx}\,dx$
-$\qquad=\dfrac{ab}{i2n}\int_{\pi-c}^{\pi}(1-e^{-i2nx})\,dx-aq_{-2n}\int_{\pi-c}^{\pi}xe^{-i2nx}\,dx+a\pi q_{-2n}\int_{\pi-c}^{\pi}e^{-i2nx}\,dx$
-$\qquad=\dfrac{ab}{i2n}\biggl(c-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{aq_{-2n}}{i2n}\biggl(\pi-(\pi-c)e^{i2nc}-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{a\pi q_{-2n}}{i2n}(e^{i2nc}-1).$
-Therefore,
-$D_{1}(n,p)=\frac{i}{4\pi n}\int_{0}^{\pi}p(x)Q(x,n)e^{-i2nx}\,dx-\frac{iQ_{n,0}}{4\pi n}\int_{0}^{\pi}p(x)e^{-i2nx}\,dx$
-$\qquad=\frac{i}{4\pi n}\biggl(\frac{ab\pi}{i2n}+\frac{(a-b)(e^{i2nc}-1)q_{-2n}}{4n^{2}}-\frac{a\pi(e^{i2nc}-1)q_{-2n}}{i2n}\biggr)-\frac{iQ_{n,0}p_{2n}}{4n}$
-and substituting $p_{2n}=q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$ in the last expression, we have
-$D_{1}(n,p)=\frac{ab}{8n^{2}}-\frac{i(b-a)(b+3a)(e^{i2nc}-1)^{2}}{32\pi n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{16\pi^{2}n^{4}}.$ (41)
-Now let us calculate the terms in $D_{2}(n,p)$:
-$\int_{0}^{\pi-c}p(x)G(x,n)e^{-i2nx}\,dx=\int_{0}^{\pi-c}b\biggl(\dfrac{a}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)\biggr)e^{-i2nx}\,dx$
-$\qquad=\dfrac{ab}{i2n}\int_{0}^{\pi-c}(1-e^{-i2nx})\,dx+b\dfrac{\pi}{2}q_{-2n}\int_{0}^{\pi-c}xe^{-i(2n-1)x}\,dx-b\dfrac{\pi}{2}q_{-2n}\int_{0}^{\pi-c}xe^{-i2nx}\,dx$
-$\qquad=\dfrac{ab}{i2n}\biggl(\pi-c+\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{b\pi q_{-2n}}{i2(2n-1)}\biggl((\pi-c)e^{i(2n-1)c}+\dfrac{e^{i(2n-1)c}+1}{i(2n-1)}\biggr)$
-$\qquad+\dfrac{b\pi q_{-2n}}{i4n}\biggl((\pi-c)e^{i2nc}+\dfrac{e^{i2nc}-1}{i2n}\biggr)$
-and
-$\int_{\pi-c}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx=\int_{\pi-c}^{\pi}a\biggl(\dfrac{b}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)+\pi q_{-2n}\biggr)e^{-i2nx}\,dx$
-$\qquad=\dfrac{ab}{i2n}\int_{\pi-c}^{\pi}(1-e^{-i2nx})\,dx+a\dfrac{\pi}{2}q_{-2n}\int_{\pi-c}^{\pi}xe^{-i(2n-1)x}\,dx-a\dfrac{\pi}{2}q_{-2n}\int_{\pi-c}^{\pi}xe^{-i2nx}\,dx+a\pi q_{-2n}\int_{\pi-c}^{\pi}e^{-i2nx}\,dx$
-$\qquad=\dfrac{ab}{i2n}\biggl(c-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{a\pi q_{-2n}}{i2(2n-1)}\biggl(\pi-(\pi-c)e^{i(2n-1)c}+\dfrac{-e^{i(2n-1)c}+1}{i(2n-1)}\biggr)$
-$\qquad+\dfrac{a\pi q_{-2n}}{i4n}\biggl(\pi-(\pi-c)e^{i2nc}-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{a\pi q_{-2n}}{i2n}(e^{i2nc}-1)$
-give
-$D_{2}(n,p)=\frac{i}{4\pi n}\int_{0}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx$
-$\qquad=\frac{i}{4\pi n}\biggl(\frac{ab\pi}{i2n}+\frac{a\pi^{2}(1-e^{i(2n-1)c})q_{-2n}}{i2(2n-1)}-\frac{\pi(b+a+(b-a)e^{i(2n-1)c})q_{-2n}}{2(2n-1)^{2}}$
-$\qquad+\frac{a\pi^{2}(1-e^{i2nc})q_{-2n}}{i4n}+\frac{\pi(b-a)(1-e^{i2nc})q_{-2n}}{8n^{2}}+\dfrac{a\pi(e^{i2nc}-1)q_{-2n}}{i2n}\biggr).$
-Using $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$, we obtain
 $D_{2}(n,p)=\frac{ab}{8n^{2}}+\frac{i(b-a)(1-e^{i(2n-1)c})(e^{i2nc}-1)}{16n^{2}(2n-1)}+\frac{(b^{2}-a^{2}+(b-a)^{2}e^{i(2n-1)c})(e^{i2nc}-1)}{16\pi n^{2}(2n-1)^{2}}$
 $\qquad-\frac{ia(b-a)(e^{i2nc}-1)^{2}}{32n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{64\pi n^{4}}+\frac{ia(b-a)(e^{i2nc}-1)^{2}}{16\pi n^{3}}.$ (39)
-Thus, by (41), (42), and Lemma 1, we write
-$D(n,p)=D_{1}(n,p)+D_{2}(n,p)=\frac{ab}{4n^{2}}-\frac{i(b-a)(b+3a)(e^{i2nc}-1)^{2}}{32\pi n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{16\pi^{2}n^{4}}$
-$\qquad+\frac{i(b-a)(1-e^{i(2n-1)c})(e^{i2nc}-1)}{16n^{2}(2n-1)}+\frac{(b^{2}-a^{2}+(b-a)^{2}e^{i(2n-1)c})(e^{i2nc}-1)}{16\pi n^{2}(2n-1)^{2}}$
-$\qquad-\frac{ia(b-a)(e^{i2nc}-1)^{2}}{32n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{64\pi n^{4}}+\frac{ia(b-a)(e^{i2nc}-1)^{2}}{16\pi n^{3}}.$
-and simplifying a bit
+Thus, using (38), (39), Lemma 1, and simplifying a bit, we write
 $D(n,p)=D_{1}(n,p)+D_{2}(n,p)=\frac{ab}{4n^{2}}+\frac{i(a^{2}-b^{2})(e^{i2nc}-1)^{2}}{32\pi n^{3}}$
@@
 $\qquad+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{64\pi n^{4}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{16\pi^{2}n^{4}}+\frac{(b^{2}-a^{2}+(b-a)^{2}e^{i(2n-1)c})(e^{i2nc}-1)}{16\pi n^{2}(2n-1)^{2}}.$ (40)
-Now we calculate $B_{1,n}(n^{2})$ defined by (22) in Lemma 2 . Expressing $Q(x)$ defined in Lemma 2 for the Kronig-Penney potential as
-$Q(x)=\begin{cases}ax&\text{if }x\in[0,c],\\[5.69054pt] bx-b\pi&\text{if }x\in(c,\pi]\end{cases},$
+Now, we calculate $B_{1,n}(n^{2})$ defined by (21). Expressing $Q(x)$ defined in Lemma 2 for the Kronig-Penney potential as
+$Q(x)=\begin{cases}ax&\text{if }x\in[0,c]\\[5.69054pt] bx-b\pi&\text{if }x\in(c,\pi],\end{cases}$
 by direct calculation, we obtain
 $B_{1,n}(n^{2})=-\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx=\frac{ab\cos 2nc}{2n^{2}}-\frac{(b^{2}-a^{2})\sin 2nc}{4\pi n^{3}}.$ (41)
-Therefore, by
-$C_{k}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos kxdx=\frac{a-b}{\pi k}\sin kc,$
-(40), (43), and (44), we can calculate:
-$-C_{2n}-\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)$
-$\qquad=-\frac{a-b}{2\pi n}\sin 2nc+\frac{ab\cos 2nc}{2n^{2}}+\frac{ab}{4n^{2}}+O\bigl(\frac{1}{n^{3}}\bigr).$
+Therefore, by $C_{k}=\dfrac{a-b}{\pi k}\sin kc$, (37), (40), and (41), we can calculate:
+$-C_{2n}+\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)$
+$\qquad=-\frac{a-b}{2\pi n}\sin 2nc-\frac{ab\cos 2nc}{2n^{2}}+\frac{ab}{4n^{2}}+O\bigl(\frac{1}{n^{3}}\bigr).$
 Thus, by (30), we have the following result:
@@
 If $q(x)$ has the form (9), then the following asymptotic formula holds:
-$\lambda_{n}=n^{2}+\frac{b-a}{2\pi n}\sin 2nc+\frac{ab(2\cos 2nc+1)}{4n^{2}}+O\bigl(\frac{1}{n^{3}}\bigr).$
+$\lambda_{n}=n^{2}+\frac{b-a}{2\pi n}\sin 2nc+\frac{ab(1-2\cos 2nc)}{4n^{2}}+O\bigl(\frac{1}{n^{3}}\bigr).$
 3 Estimates for the Small Eigenvalues
@@
 $\lambda-n^{2}+C_{2n}-\sum\limits_{k=1}^{\infty}a_{k}(\lambda)=0,$ (42)
-in the set $I_{n}:=[n^{2}-M,n^{2}+M]$, where $M=\max\{|a|,b\}$, $C_{k}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos kxdx$ and the infinite series $a_{k}$ is defined by (5), if the following condition holds:
+in the set $I_{n}:=[n^{2}-M,n^{2}+M]$, where $M=\max\{|a|,b\}$, and the infinite series $a_{k}$ is defined by (5), if the following condition holds:
 Condition 1
-To consider the $n$th eigenvalue $\lambda_{n}$ for $n>1,$ we assume that the condition M<(2n1)/2M<(2n-1)/2 holds. In case $n=1$, we assume that the condition $M\leq 1/2$ holds.
-Then, to use numerical methods we take finite sums instead of the infinite series in the equations obtained.
-To approximate the roots of the equation (45), we use the fixed point iteration.
-It can also be used the Newton-Raphson method but in this case it is necessary to compute the derivative of the function $K_{n}(\lambda)$ defined by (55).
-Then, using the Banach fixed point theorem, we prove that each of these equations containing the finite sums has a unique solution in the appropriate set $I_{n}$ (see Theorem 4).
+To consider the $n$th eigenvalue $\lambda_{n}$, for $n>1$, we assume that the condition M<(2n1)/2M<(2n-1)/2 holds. In case $n=1$, we assume that the condition $M\leq 1/2$ holds.
+Then, to use numerical methods, we take finite sums instead of the infinite series in (42).
+To approximate the roots of equation (42), we use the fixed point iteration.
+Using the Banach fixed point theorem, we prove that the equation containing the finite sums has a unique solution in the appropriate set $I_{n}$ (see Theorem 4).
 Moreover, we give error analysis (see Theorem 4 and Theorem 5) and present a numerical example.
 Now, we state some preliminary facts.
-It is well known that the spectrum of the operator $L(q)$ is discrete and for large enough $n$, there is one eigenvalue (counting with multiplicity) in the neighborhood of $n^{2}$.
+It is well known that the spectrum of the operator $L(q)$ is discrete and for large enough $n$, there is one eigenvalue (counting multiplicity) in the neighborhood of $n^{2}$.
 See the basic and detailed classical results in [1, 2, 4, 5] and references therein.
-The eigenvalues of the operators $L(0)$ are $n^{2}$, for $n\in\mathbb{Z}^{+}$ and all the eigenvalues of $L(0)$ are simple.
-It is also known that [6]
-$|\lambda_{n}-n^{2}|\leq M,$
-for $n\geq 1$. Therefore, we have
+The eigenvalues of the operator $L(0)$ are $n^{2}$, for $n\in\mathbb{N}$, where $\mathbb{N}$ is the set of positive integers, and all the eigenvalues of $L(0)$ are simple.
+It is also known that [6], $|\lambda_{n}-n^{2}|\leq M$, for $n\geq 1$. Therefore, we have
 $n^{2}-M\leq\lambda_{n}\leq n^{2}+M,$ (43)
-for $n\geq 1$. If $k\neq\pm n$, then
-$|\lambda_{n}-k^{2}|\geq|n^{2}-k^{2}|-M=|n-k||n+k|-M$
-$\qquad\geq(2n-1)-M,$ (47)
-for $n\geq 1$ and under the assumptions about $M$ given in Condition 1. In particular, if $n=1$, we have $|\lambda_{1}|\leq 1+M$ and
-$|\lambda_{1}-k^{2}|\geq||\lambda_{1}|-k^{2}|\geq 4-|\lambda_{1}|\geq 3-M,$
-for $|k|\geq 2$. Besides, if $n\geq 2$, we have $|\lambda_{n}|\geq|\lambda_{2}|\geq 4-M$ and
-$|\lambda_{n}-k^{2}|\geq||\lambda_{2}|-k^{2}|\geq|\lambda_{2}|-1\geq 3-M,$
-for $k\neq\pm n$.
+for $n\geq 1$.
+If $k\neq\pm n$, then $|\lambda_{n}-k^{2}|\geq|n^{2}-k^{2}|-M=|n-k||n+k|-M\geq(2n-1)-M$, for $n\geq 1$ and under the assumptions about $M$ given in Condition 1.
+In particular, if $n=1$, we have $|\lambda_{1}|\leq 1+M$ and $|\lambda_{1}-k^{2}|\geq||\lambda_{1}|-k^{2}|\geq 4-|\lambda_{1}|\geq 3-M$, for $|k|\geq 2$.
+Besides, if $n\geq 2$, we have $|\lambda_{n}|\geq|\lambda_{2}|\geq 4-M$ and $|\lambda_{n}-k^{2}|\geq||\lambda_{2}|-k^{2}|\geq|\lambda_{2}|-1\geq 3-M$, for $k\neq\pm n$.
 We stress that, the iteration formula (2) was used in [13] for large eigenvalues to obtain asymptotic formulas.
@@
 We also note that, it is not easy to give such conditions, there are many technical calculations.
-We remind that, since the potential $q$ is of the form (9), we have
-$C_{k}=\frac{a-b}{\pi k}\sin kc,\qquad C_{-k}=C_{k},$
-for $k=1,2,\ldots$.
+We remind that, since the potential $q$ is of the form (9), we have $C_{k}=\dfrac{a-b}{\pi k}\sin kc$ and $C_{-k}=C_{k}$, for $k=1,2,\ldots$.
 Now, in order to give the main results, we prove the following lemma.
@@
 (a) Since $\Psi_{n}||=1$ and $\sin kx||=\sqrt{\pi}/\sqrt{2}$, by the Schwarz inequality, we have $|(q\Psi_{n},\sin kx)|\leq M\sqrt{\pi}/\sqrt{2}$.
-Considering the number of terms and the greatest summands of $R_{2m}(\lambda)$ in absolute value, by (46) and (47) we obtain,
+Considering the number of terms and the greatest summands of $R_{2m}(\lambda)$ in absolute value, by (43), we obtain,
 $|R_{2m}(\lambda_{1})|<\frac{4^{m}|C_{1}||C_{2}|^{2m}M\sqrt{\pi}/\sqrt{2}}{(2n-1-M)^{2m+1}}<\frac{(b-a)}{\sqrt{2\pi}}(\frac{2}{\pi})^{2m}.$
 Therefore, $\lim_{m\rightarrow\infty}R_{m}(\lambda)=0$.
-(b) Suppose the contrary, $(\Psi_{n},\sin nx)=0$. Since the system of root functions
-$\{\sqrt{2}\sin kx/\sqrt{\pi}:k\in\mathbb{Z}^{+}\}$
-of $L(0)$ forms an orthonormal basis for $L_{2}[0,\pi]$, we have the decomposition
-$\dfrac{\pi}{2}\Psi_{n}=(\Psi_{n},\sin nx)\sin nx+\sum_{k\in\mathbb{Z}^{+},k\neq n}\left(\Psi_{n},\sin kx\right)\sin kx$
+(b) Suppose the contrary, $(\Psi_{n},\sin nx)=0$.
+Since the system of root functions $\{\sqrt{2}\sin kx/\sqrt{\pi}:k\in\mathbb{N}\}$ of $L(0)$ forms an orthonormal basis for $L_{2}[0,\pi]$, we have the decomposition
+$\dfrac{\pi}{2}\Psi_{n}=(\Psi_{n},\sin nx)\sin nx+\sum_{k\in\mathbb{N},k\neq n}\left(\Psi_{n},\sin kx\right)\sin kx$
 for the normalized eigenfunction $\Psi_{n}$ corresponding to the eigenvalue $\lambda_{n}$ of $L(q)$. By Parseval’s equality, we obtain
-$\sum_{k\in\mathbb{Z}^{+},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\frac{\pi}{2}.$
-Using the relation,
-$(\lambda_{N}-n^{2})(\Psi_{N},\sin nx)=(q\Psi_{N},\sin nx)$
-which is obtained from
-$\Psi_{N}^{\prime\prime}(x)+q(x)\Psi_{N}(x)=\lambda_{N}\Psi_{N}(x),$
-by multiplying both sides of the equality by $\sin nx$, where $\Psi_{N}(x)$ is an eigenfunction corresponding to the eigenvalue $\lambda_{N}$, the Bessel inequality and (47), we obtain
-$\sum_{k\in\mathbb{Z}^{+},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\sum_{k\in\mathbb{Z}^{+},k\neq n}\frac{|(q\Psi_{n},\sin kx)|^{2}}{|\lambda_{n}-k^{2}|^{2}}$
-$\qquad\leq\frac{1}{(2n-1-M)^{2}}\sum_{k\in\mathbb{Z},k\neq\pm n}|(q\Psi_{n},\sin kx)|^{2}<\frac{\pi(M)^{2}}{2(2n-1-M)^{2}}<\frac{\pi}{2},$
-which contradicts $\sum_{k\in\mathbb{Z}^{+},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\pi/2$ and completes the proof.
-Before stating the main results, we want to approximate the following series defined in Lemma 1 in a different way.
-$A_{1,n}(n^{2})$ $=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}{}^{2}}{n^{2}-(n+k)^{2}}=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}^{2}}{(2n+k)(-k)}=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq-2n\end{subarray}}^{-1}\frac{C_{k}^{2}}{(2n+k)(-k)}+\sum_{k=1}^{\infty}\frac{C_{k}^{2}}{(2n+k)(-k)}$
-$=\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\biggl(\frac{C_{k}^{2}}{k(2n-k)}-\frac{C_{k}^{2}}{k(2n+k)}\biggr)=\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{2C_{k}^{2}}{(2n-k)(2n+k)}.$ (48)
-To this end, we introduce the integral
-$I=\int_{0}^{\pi}(P(x,n)-\frac{C_{4n}}{2n}\sin 2nx)^{2}\cos 4nxdx,$ (49)
-where
-$P(x,n)=\int_{0}^{x}q(t)\cos 2ntdt-C_{2n}x.$
-It is obvious that
-$P(0,n)=P(\pi,n)=0,\qquad P^{\prime}(x,n)=q(x)\cos 2nx-C_{2n},$
-and
-$P(x,n)=\begin{cases}\dfrac{a}{2n}\sin 2nx-C_{2n}x&\text{if }x\in[0,c],\\[5.69054pt] \dfrac{b}{2n}\sin 2nx-C_{2n}x+C_{2n}\pi&\text{if }x\in(c,\pi].\end{cases}$
-The Fourier sine coefficients of $P(x,n)$ are
-$P_{k}(n)=\int_{0}^{\pi}P(x,n)\sin kxdx=\frac{\pi(C_{2n+k}+C_{2n-k})}{2k},\qquad P_{2n}(n)=\frac{\pi C_{4n}}{4n}.$
-Using the Fourier decomposition
-$P(x,n)-\frac{C_{4n}}{2n}\sin 2nx=\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{C_{2n+k}+C_{2n-k}}{k}\sin kx$
-of $P(x,n)-\dfrac{C_{4n}}{2n}\sin 2nx$ in the integral $I$ defined by (49), we obtain
-$I$ $=\int_{0}^{\pi}(P(x,n)-\frac{C_{4n}}{2n}\sin 2nx)^{2}\cos 4nxdx=\int_{0}^{\pi}\biggl(\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{C_{2n+k}+C_{2n-k}}{k}\sin kx\biggr)^{2}\cos 4nxdx$
-$=\int_{0}^{\pi}\biggl(\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{C_{4n-k}+C_{k}}{2n-k}\sin(2n-k)x\frac{C_{4n+k}+C_{k}}{2n+k}\sin(2n+k)x\biggr)\cos 4nxdx$
-$=-\frac{\pi}{4}\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{(C_{4n-k}+C_{k})(C_{4n+k}+C_{k})}{(2n-k)(2n+k)}$
-and therefore
-$\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{(C_{4n-k}+C_{k})(C_{4n+k}+C_{k})}{(2n-k)(2n+k)}=\frac{-4I}{\pi}.$ (50)
-Now, the integration by parts in the integral $I$ defined by (49) gives
-$I$ $=-\frac{1}{2n}\int_{0}^{\pi}P(x,n)(q(x)\cos 2nx-C_{2n})\sin 4nxdx+\frac{C_{4n}}{4n^{2}}\int_{0}^{\pi}\sin 2nx(q(x)\cos 2nx-C_{2n})\sin 4nxdx$
-$=-\frac{1}{4n}\int_{0}^{\pi}P(x,n)q(x)[\sin 6nx+\sin 2nx]dx+\frac{C_{2n}}{2n}\int_{0}^{\pi}P(x,n)\sin 4nxdx+\frac{C_{4n}}{8n^{2}}\int_{0}^{\pi}q(x)[1-\cos 8nx]dx$
-$=-\frac{1}{4n}I_{1}+\frac{C_{2n}}{2n}I_{2}-\frac{\pi C_{4n}C_{8n}}{8n^{2}},$
-where
-$I_{1}=\int_{0}^{\pi}P(x,n)q(x)[\sin 6nx+\sin 2nx]dx,\qquad I_{2}=\int_{0}^{\pi}P(x,n)\sin 4nxdx.$
-By direct calculations, we obtain
-$I_{1}$ $=-\frac{ab\pi}{4n}+\frac{(b^{2}-a^{2})\sin 8nc}{32n^{2}}-\frac{(b-a)^{2}}{16\pi n^{3}}+\frac{(b-a)^{2}\cos 4nc}{18\pi n^{3}}+\frac{(b-a)^{2}\cos 8nc}{144\pi n^{3}},$
-$I_{2}$ $=\frac{(b-a)\sin 2nc}{16n^{2}}-\frac{(b-a)\sin 6nc}{48n^{2}},$
-and
-$I$ $=\frac{ab\pi}{16n^{2}}-\frac{(b^{2}-a^{2})\sin 8nc}{128n^{3}}+\frac{13(b-a)^{2}\cos 4nc}{3.2^{9}\pi n^{4}}$
-$+\frac{(b-a)^{2}}{128\pi n^{4}}-\frac{5(b-a)^{2}\cos 8nc}{9.2^{7}\pi n^{4}}+\frac{(b-a)^{2}\cos 12nc}{2^{9}\pi n^{4}}.$
-Therefore, by (48)-(50) we have
-$|A_{1,n}(n^{2})|$ $<\frac{12}{\pi}|I|=\frac{3ab}{4n^{2}}-\frac{3(b^{2}-a^{2})\sin 8nc}{32\pi n^{3}}$ (51)
-$+\frac{13(b-a)^{2}\cos 4nc}{2^{7}\pi^{2}n^{4}}+\frac{3(b-a)^{2}}{32\pi^{2}n^{4}}-\frac{5(b-a)^{2}\cos 8nc}{96\pi^{2}n^{4}}+\frac{3(b-a)^{2}\cos 12nc}{2^{7}\pi^{2}n^{4}}.$
-Now, letting $m$ tend to infinity in equation (2), we obtain the following main results.
+$\sum_{k\in\mathbb{N},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\frac{\pi}{2}$ (44)
+On the other hand, using the relation $(\lambda_{N}-n^{2})(\Psi_{N},\sin nx)=(q\Psi_{N},\sin nx)$, the Bessel inequality and (43), we obtain
+$\sum_{k\in\mathbb{N},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\sum_{k\in\mathbb{N},k\neq n}\frac{|(q\Psi_{n},\sin kx)|^{2}}{|\lambda_{n}-k^{2}|^{2}}$
+$\qquad\leq\frac{1}{(2n-1-M)^{2}}\sum_{k\in\mathbb{Z},k\neq\pm n}|(q\Psi_{n},\sin kx)|^{2}<\frac{\pi(M)^{2}}{2(2n-1-M)^{2}}<\frac{\pi}{2}.$
+which contradicts Parseval’s equality (44).
+Before giving the main results, we calculate the number $C_{2n}+A_{1,n}(n^{2})-B_{1,n}(n^{2})$ by (14), (37), (40), and (41):
+$-C_{2n}+A_{1,n}(n^{2})-B_{1,n}(n^{2})=\frac{b-a}{2\pi n}\sin 2nc+\frac{ab(1-2\cos 2nc)}{4n^{2}}+\frac{(b-a)(\cos 2nc-1)(2a\sin 2nc+ib)}{16n^{3}}$
+$\qquad+\frac{(b^{2}-a^{2})(2\sin 2nc(2+\cos 2nc)+i(\cos 2nc-1))}{16\pi n^{3}}+\frac{(b-a)^{2}\cos 2nc(\cos 2nc-1)}{16\pi^{2}n^{4}}$
+$\qquad-\frac{-(b-a)[2(\sin 2nc+\sin(2n-1)c-\sin(4n-1)c)+ib(1-e^{-i2nc})(e^{i2nc}+e^{ic})]}{16n^{2}(2n-1)}$
+$\qquad+\frac{(b-a)^{2}(\cos 4nc-\cos 2nc)}{32\pi n^{4}}+\frac{(b^{2}-a^{2})(e^{-i2nc}-1)-(b-a)^{2}(e^{ic}+e^{i(2n-1)c}-2\cos(4n-1)c)}{16\pi n^{2}(2n-1)^{2}}.$ (45)
+Now, we state the following main result.
 Theorem 3
-If the assumptions about $M$ given in Condition 1 hold, then $\lambda_{n}$ is an eigenvalue of $L(q)$ if and only if it is the root of the equation
-$\lambda-n^{2}+C_{2n}-\sum\limits_{k=1}^{\infty}a_{k}(\lambda)=0$ (52)
-in the set $I_{n}:=[n^{2}-M,n^{2}+M]$.
+If the assumptions about $M$ given in Condition 1 hold, then $\lambda_{n}$ is an eigenvalue of $L(q)$ if and only if it is the root of equation (42) in the set $I_{n}:=[n^{2}-M,n^{2}+M]$.
 Proof. (a) By Lemma 4, letting $m$ tend to infinity in equation (2), we obtain
-$\lambda_{n}-n^{2}+C_{2n}-\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})=0,$
-where $C_{k}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos kxdx$.
-Now, we prove that the root of (52) lying in the interval $I_{n}$ is an eigenvalue of the operator $L$. To do this, we first give the following relations:
+$\lambda_{n}-n^{2}+C_{2n}-\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})=0.$
+Now, we prove that the root of (42) lying in the interval $I_{n}$ is an eigenvalue of the operator $L$. To do this, we first give the following relation:
 $a_{1}(\lambda)=a_{1}(n^{2})+\sum_{\begin{subarray}{c}n_{1}=-\infty\\ n_{1}\neq 0,2n\end{subarray}}^{\infty}\frac{C_{n_{1}}(C_{n_{1}}-C_{2n+n_{1}})(\lambda-n^{2})}{n_{1}(2n+n_{1})[\lambda-(n+n_{1})^{2}]}$
-and
-$a_{1}(n^{2})=A_{1,n}(n^{2})+B_{1,n}(n^{2}),$ (53)
-where $A_{1,n}(n^{2})$ and $B_{1,n}(n^{2})$ are given by (11) and (12), respectively.
+where $a_{1}(\lambda)$ is defined in (3), $a_{1}(n^{2})=A_{1,n}(n^{2})-B_{1,n}(n^{2})$, $A_{1,n}(n^{2})$ and $B_{1,n}(n^{2})$ are given in (11).
 The equation
-$F_{n}(\lambda):=\lambda-n^{2}+C_{2n}-A_{1,n}(n^{2})-B_{1,n}(n^{2})$
+$F_{n}(\lambda):=\lambda-n^{2}+C_{2n}-A_{1,n}(n^{2})+B_{1,n}(n^{2})$
 has one root in the interval $I_{n}=[n^{2}-M,n^{2}+M]$ and
-$|F_{n}(\lambda)|\geq|\lambda-n^{2}|-|C_{2n}|-|A_{1,n}(n^{2})|-|B_{1,n}(n^{2})|.$
-Estimating $|A_{1,n}(n^{2})|$ and $|B_{1,n}(n^{2})|$ by (51) and (44), we have
+$|F_{n}(\lambda)|\geq|\lambda-n^{2}|-|C_{2n}-A_{1,n}(n^{2})+B_{1,n}(n^{2})|.$
+Estimating $|C_{2n}-A_{1,n}(n^{2})+B_{1,n}(n^{2})|$ by (45), we have
 $|F_{n}(\lambda)|>0.16,$
@@
 $|a_{1}(\lambda)-a_{1}(n^{2})|<\frac{16}{45\pi^{2}},\qquad|a_{k}(\lambda)|<\frac{\sqrt{2}^{k}}{\pi^{k+1}},$
-for $k\geq 2$, where $a_{1}(n^{2})$ is defined by (53). Therefore, it follows by the geometric series formula that
+for $k\geq 2$. Therefore, it follows by the geometric series formula that
 $|a_{1}(\lambda)-a_{1}(n^{2})|+\sum\limits_{k=2}^{\infty}|a_{k}(\lambda)|<\frac{16}{45\pi^{2}}+\frac{\pi+1}{\pi^{2}(\pi-1)}<0.1534$
@@
 $|G_{n}(\lambda)-F_{n}(\lambda)|=\biggl|a_{1}(\lambda)-a_{1}(n^{2})+\sum\limits_{k=2}^{\infty}a_{k}(\lambda)\biggr|$
-$\qquad\leq|a_{1}(\lambda)-a_{1}(n^{2})|+\sum\limits_{k=2}^{\infty}|a_{k}(\lambda)|$
-$\qquad<\frac{16}{45\pi^{2}}+\frac{\pi+1}{\pi^{2}(\pi-1)}<0.1534$
+$\qquad\leq|a_{1}(\lambda)-a_{1}(n^{2})|+\sum\limits_{k=2}^{\infty}|a_{k}(\lambda)|<\frac{16}{45\pi^{2}}+\frac{\pi+1}{\pi^{2}(\pi-1)}<0.1534$
 for all $\lambda$ from the boundary of $I_{n}$, for $n=1$.
@@
 Thus, $\lambda\in I_{n}$ is an eigenvalue of $L(q)$ if and only if, it is the root of (42).
-Now, we can use numerical methods by taking finite sums instead of the infinite series in (52) and obtain
+Now, we can use numerical methods by taking finite sums instead of the infinite series in (42) and write
 $\lambda-n^{2}+C_{2n}-\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda)=0,$
-where
-$a_{r,k,n}(\lambda)=\sum_{n_{1},n_{2},...,n_{k}=-r}^{r}\frac{C_{n_{1}}C_{n_{2}}\cdots C_{n_{k}}(C_{n_{1}+n_{2}+\cdots+n_{k}}-C_{n_{1}+n_{2}+\cdots+n_{k}+2n})}{[\lambda-(n+n_{1})^{2}]\cdots[\lambda-(n+n_{1}+\cdots+n_{k})^{2}]}.$ (54)
-Define the function
-$K_{n}(\lambda):=\lambda-n^{2}-g_{n}(\lambda),$ (55)
-where
-$g_{n}(\lambda)=-C_{2n}+\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda).$ (56)
+where $a_{r,k,n}(\lambda)$ is obtained from (5) by taking finite sums instead of the infinite series, namely,
+$a_{r,k,n}(\lambda):=\sum_{n_{1},n_{2},...,n_{k}=-r}^{r}\frac{C_{n_{1}}C_{n_{2}}\cdots C_{n_{k}}(C_{n_{1}+n_{2}+\cdots+n_{k}}-C_{n_{1}+n_{2}+\cdots+n_{k}+2n})}{[\lambda-(n+n_{1})^{2}]\cdots[\lambda-(n+n_{1}+\cdots+n_{k})^{2}]}.$
+Define the functions
+$K_{n}(\lambda):=\lambda-n^{2}-g_{n}(\lambda),\qquad g_{n}(\lambda)=-C_{2n}+\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda).$ (46)
 Then,
@@
 $\qquad+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M](1-L_{n})}.$ (49)
-Proof. First we prove (58) by using the mean-value theorem. To do this, we estimate $|g_{n}^{\prime}(\lambda)|=|\dfrac{d}{d\lambda}g_{n}(\lambda)|$. By (56), we have
+Proof. First, we prove (48) by using the mean-value theorem. To do this, we estimate $|g_{n}^{\prime}(\lambda)|=|\dfrac{d}{d\lambda}g_{n}(\lambda)|$. By (46), we have
 $|g_{n}^{\prime}(\lambda)|=\biggl|\sum\limits_{k=1}^{s}\frac{d}{d\lambda}a_{r,k,n}(\lambda)\biggr|\leq\sum\limits_{k=1}^{s}\bigl|\frac{d}{d\lambda}a_{r,k,n}(\lambda)\bigr|.$
-First let us estimate the summands of the term $|\dfrac{d}{d\lambda}a_{r,k,n}(\lambda)|$:
+By estimating the summands of the term $|\dfrac{d}{d\lambda}a_{r,k,n}(\lambda)|$, we write
 $\biggl|\frac{d}{d\lambda}(a_{r,k,n}(\lambda))\biggr|=\biggl|\sum_{n_{1},n_{2},...,n_{k}=-r}^{r}\frac{d}{d\lambda}\frac{C_{n_{1}}C_{n_{2}}\cdots C_{n_{k}}(C_{n_{1}+n_{2}+\cdots+n_{k}}-C_{n_{1}+n_{2}+\cdots+n_{k}+2n})}{[\lambda-(n+n_{1})^{2}]\cdots[\lambda-(n+n_{1}+\cdots+n_{k})^{2}]}\biggr|$
@@
 $\qquad\leq\biggl|\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})-\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|.$ (52)
-For the estimation of the right-hand side of (62), we obtain
+For the estimation of the right-hand side of (52), we have
 $\biggl|\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})-\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|$
@@
 $\biggl|\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})-\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|<\sum\limits_{k=s+1}^{\infty}\frac{(b-a)^{k+1}}{2M\sqrt{2}^{k}\pi^{k+1}(2n-1-M)^{k}}+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M]}$
-$\qquad=\frac{(b-a)^{s+2}}{2M\sqrt{2}^{s}\pi^{s+1}(2n-1-M)^{s}[\sqrt{2}\pi(2n-1-M)-(b-a)]}+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M]}$ (63)
-Thus, by (62) and (63), we obtain
+$\qquad=\frac{(b-a)^{s+2}}{2M\sqrt{2}^{s}\pi^{s+1}(2n-1-M)^{s}[\sqrt{2}\pi(2n-1-M)-(b-a)]}+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M]}.$ (53)
+Thus, by (52) and (53), we have
 $|K_{n}(\lambda_{n})-K_{n}(\rho_{n})|<\frac{(b-a)^{s+2}}{2M\sqrt{2}^{s}\pi^{s+1}(2n-1-M)^{s}[\sqrt{2}\pi(2n-1-M)-(b-a)]}$
@@
 $x_{n,i+1}=n^{2}+g_{n}(x_{n,i}),$ (56)
-where $g_{n}(x)$ is defined by (56). Since
-$|g_{n}(\lambda_{n})|=\biggl|-C_{2n}+\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|$
-$\qquad\leq|C_{2n}|+\sum\limits_{k=1}^{s}|a_{r,k,n}(\lambda_{n})|,$ (67)
+where $g_{n}(x)$ is defined in (46). Since
+$|g_{n}(\lambda_{n})|=\biggl|-C_{2n}+\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|\leq|C_{2n}|+\sum\limits_{k=1}^{s}|a_{r,k,n}(\lambda_{n})|,$ (57)
 first let us estimate $|a_{r,k,n}(\lambda)|$:
@@
 On the other hand, writing (2n1)(2n-1) instead of (2n1)M(2n-1)-M in (57) and (58), we obtain
-$|g_{n}(n^{2})|\leq|C_{2n}|+\sum\limits_{k=1}^{s}|a_{r,k,n}(n^{2})|$
-$\qquad<\frac{(b-a)}{2\pi n}+\frac{(b-a)^{2}}{M\pi^{2}(2n-1)}+\sum\limits_{k=2}^{r}\frac{(b-a)^{k+1}}{2M\sqrt{2}^{k}\pi^{k+1}(2n-1)^{k}}$
+$|g_{n}(n^{2})|\leq|C_{2n}|+\sum\limits_{k=1}^{s}|a_{r,k,n}(n^{2})|<\frac{(b-a)}{2\pi n}+\frac{(b-a)^{2}}{M\pi^{2}(2n-1)}+\sum\limits_{k=2}^{r}\frac{(b-a)^{k+1}}{2M\sqrt{2}^{k}\pi^{k+1}(2n-1)^{k}}$
 $\qquad<\frac{(b-a)}{2\pi n}+\frac{(b-a)^{2}}{M\pi^{2}(2n-1)}+\frac{(b-a)^{3}}{2\sqrt{2}M\pi^{2}(2n-1)[\sqrt{2}\pi(2n-1)-(b-a)]},$ (59)
@@
 for $i=1,2,3,\ldots$, where $L_{n}$ is defined in (48) in Theorem 4.
-Proof. Without loss of generality we can take $x_{n,0}=n^{2}$. By (57), (58), and (66) we have
+Proof. Without loss of generality, we can take $x_{n,0}=n^{2}$. By (47), (48), and (56), we have
 $|x_{n,i}-\rho_{n}|=|n^{2}+g_{n}(x_{n,i-1})-(n^{2}+g_{n}(\rho_{n}))|$
 $\qquad=|g_{n}(x_{n,i-1})-g_{n}(\rho_{n})|<L_{n}|x_{n,i-1}-\rho_{n}|<(L_{n})^{i}|x_{n,0}-\rho_{n}|.$ (61)
-Therefore it is enough to estimate $|x_{n,0}-\rho_{n}|$. By definitions of $\rho_{n}$ and $x_{n,0}$ we obtain
+Therefore, it is enough to estimate $|x_{n,0}-\rho_{n}|$. By definitions of $\rho_{n}$ and $x_{n,0}$, we obtain
 $\rho_{n}-x_{n,0}=g_{n}(\rho_{n})+n^{2}-x_{n,0}=g_{n}(\rho_{n})-g_{n}(x_{n,0})+g_{n}(n^{2})$
-and by the mean value theorem there exists $x\in[n^{2}-M,n^{2}+M]$ such that
+and by the mean value theorem, there exists $x\in[n^{2}-M,n^{2}+M]$ such that
 $g_{n}(\rho_{n})-g_{n}(x_{n,0})=g_{n}^{\prime}(x)(\rho_{n}-x_{n,0}).$
@@
 In our calculations, we take $s=r=5$.
-$\lambda_{1}$ $=0.984205232093$
-$\lambda_{2}$ $=4.003110832419$
-$\lambda_{3}$ $=9.013023482675$
-$\lambda_{4}$ $=16.018415152245$
-$\lambda_{5}$ $=25.010308870985$
-$\lambda_{6}$ $=36.010794654577.$
+$\lambda_{1}$ $=0.984205232093,\quad\lambda_{2}=4.003110832419,\quad\lambda_{3}=9.013023482675$
+$\lambda_{4}$ $=16.018415152245,\quad\lambda_{5}=25.010308870985,\quad\lambda_{6}=36.010794654577.$
 Usually it takes 8108-10 iterations with the tolerance 1e181e-18 by the fixed point iteration method, even if we choose an initial value that is not too close to the exact value, which means that convergence is quite fast.
-Declarations
-  • Funding
-    Not applicable.
-  • Conflict of interest
-    The author states that there is no conflict of interest.
-  • Ethics approval
-    Not applicable.
-  • Consent to participate
-    Not applicable.
-  • Consent for publication
-    Not applicable.
-  • Availability of data and materials
-    Not applicable.